### v2

# On finite group scheme-theoretical categories, II

Abstract.
Let ${\mathscr{C}}:={\mathscr{C}}(G,\omega,H,\psi)$ be a finite group scheme-theoretical category over an algebraically closed field of characteristic $p\geq 0$ [G1]. For any indecomposable exact module category over $\mathscr{C}$}, we classify its simple objects and provide an expression for their projective covers in terms of double cosets and projective representations of certain closed subgroup schemes of $G$. This upgrades a result of Ostrik [O] for group-theoretical fusion categories in characteristic 0, and generalizes our previous work [GS] for the case $\omega=1$. As a byproduct, we describe the simples and indecomposable projectives of $\mathscr{C}$}. Finally, we apply our results to describe the blocks of the center of ${\rm Coh}(G,\omega)$.
Key words and phrases:
finite group schemes; (bi)equivariant sheaves; finite tensor categories; exact module categories
2020 Mathematics Subject Classification:
18M20, 16T05, 17B37
1. Introduction
The main purpose of this paper is to extend our previous work [GS] to arbitrary finite group scheme-theoretical categories over an algebraically closed field $k$ of characteristic $p\geq 0$ [G1]. Namely, we study the structure of indecomposable exact module categories over finite group scheme-theoretical categories ${\mathscr{C}}(G,\omega,H,\psi)$ in the presence of a not necessarily trivial $3$-cocycle $\omega$. As in [GS], we focus on classifying the simple objects of ${\mathscr{C}}(G,\omega,H,\psi)$ and describing their projective covers. Our study in particular sheds some light on the structure of ${\mathscr{C}}(G,\omega,H,\psi)$ itself, and applies to the representation category of the twisted double of a finite group scheme, for which a more precise description is obtained.
Recall that group-theoretical fusion categories [O] play a fundamental role in the theory of fusion categories in characteristic 0 (see [GS, Section 1] and references therein), which suggests that finite group scheme-theoretical categories in characteristic $p>0$ play an important role in the theory of finite tensor categories, thus making them interesting tensor categories worthwhile to study.
From an abstract point of view, a finite group scheme-theoretical category is dual to a certain finite pointed tensor category with respect to an indecomposable exact module category. More explicitly, start with a finite group scheme $G$ over $k$, and a $3$-cocycle [^1] All cochains in this paper are assumed to be normalized. $\omega\in Z^{3}(G,\mathbb{G}_{m})$ (or, a Drinfeld associator $\omega\in\mathscr{O}(G)^{\otimes 3}$ for $\mathscr{O}(G)$), and consider the finite tensor category ${\rm Coh}(G,\omega)$ of sheaves on $G$, with tensor product given by convolution and associativity by $\omega$ [G1]. Recall [G1, Theorem 5.3] that indecomposable left exact ${\rm Coh}(G,\omega)$-module categories correspond to pairs $(H,\psi)$, ${\mathscr{M}}(H,\psi)\mapsfrom(H,\psi)$, where $H\subset G$ is a closed subgroup scheme, $\psi\in C^{2}(H,\mathbb{G}_{m})$ such that $d\psi=\omega_{\mid H}$, and ${\mathscr{M}}(H,\psi)$ is the category of right $(H,\psi)$-equivariant sheaves on $(G,\omega)$. Given ${\mathscr{M}}(H,\psi)$, one calls the dual of ${\rm Coh}(G,\omega)$ with respect to ${\mathscr{M}}(H,\psi)$ a finite group scheme-theoretical category, and denote it by ${\mathscr{C}}(G,\omega,H,\psi)$ [G1]. By [EO] (see [G1, Theorem 5.7]), ${\mathscr{C}}(G,\omega,H,\psi)$ is a finite tensor category, and indecomposable left exact ${\mathscr{C}}(G,\omega,H,\psi)$-module categories correspond to indecomposable left exact ${\rm Coh}(G,\omega)$-module categories, ${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\mapsfrom{\mathscr{M}}(K,\eta)$.
Unlike in characteristic 0, finite group scheme-theoretical categories and their module categories are rarely semisimple if $p>0$, and a plethora of questions arise. In [GS], we addressed some of these questions in the case $\omega=1$, and the goal of this work is to extend that treatment to arbitrary $\omega\in Z^{3}(G,\mathbb{G}_{m})$.
The paper is organized as follows. §2 is devoted to preliminaries about finite group schemes, module categories over their categories of sheaves, and (bi)equivariant sheaves on a pair $(X,\Phi)$, where $X$ is a finite scheme and $\Phi\in C^{3}(X,\mathbb{G}_{m})$ is a $3$-cochain. In particular, we adapt [Mu, Theorem 1(B), p.112] to show that the abelian category ${\mathscr{M}}(H,\psi)$ of right $(H,\psi)$-equivariant sheaves on $(G,\omega)$ is equivalent to the category ${\rm Coh}(G/H)$ of sheaves over the finite quotient scheme $G/H$ (see Theorem 2.16). We also prove in Lemma 2.28 that there is an equivalence of abelian categories
${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\simeq{\mathscr{M}}\left((H,\psi),(K,\eta)\right),$
where ${\mathscr{M}}\left((H,\psi),(K,\eta)\right)$ is the category of $\left((H,\psi),(K,\eta)\right)$-biequivariant sheaves on $(G,\omega)$ (see Definition 2.7).
In §3, we consider the following general setting. Let $\partial:A\xrightarrow{1:1}B\times C$ be a finite group scheme embedding, and $A\backslash(B\times C)$ the finite quotient scheme with respect to the left action $\mu_{A\times(B\times C)}$ (2.27).
Fix a $3$-cochain $\Phi\in C^{3}(A\backslash(B\times C),\mathbb{G}_{m})$. Suppose that $\beta\in C^{2}(B,\mathbb{G}_{m})$ and $\gamma\in C^{2}(C,\mathbb{G}_{m})$ satisfy (2.11), and assume that there exists a $2$-cochain ${\rm W}\in C^{2}(B\times C,\mathbb{G}_{m})$ such that
$\xi:=\partial^{\otimes 2}\left(\left(\beta\times\gamma^{-1}\right){\rm W}\right)\in Z^{2}(A,\mathbb{G}_{m}).$
(Note that if $\Phi=1$, this assumption is redundant.)
Finally, define the abelian categories
$\mathscr{A}:=\mathrm{Coh}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}(A\backslash(B\times C),\Phi),$
$\mathscr{B}:=\mathrm{Coh}^{\left(\left(A\times B,\xi^{-1}\times\psi\right),\left(C,\eta\right)\right)}(B\times C,\Phi)$
(see Definition 2.7), and let
${\rm Rep}(A,\xi^{-1})_{k}:={\rm Corep}_{k}(\mathscr{O}(A)_{\xi^{-1}}),$
${\rm Rep}_{k}(A,\xi):={\rm Corep}(\mathscr{O}(A)_{\xi})_{k}$
be the categories of finite dimensional left, right comodules over the twisted coalgebras $\mathscr{O}(A)_{\xi^{-1}}$, $\mathscr{O}(A)_{\xi}$, respectively. By (2.33), there is a canonical equivalence of categories
${\rm Rep}(A,\xi^{-1})_{k}\simeq{\rm Rep}_{k}(A,\xi).$
The following result is a generalization of [GS, Theorem 3.7].
Theorem 3.8.
There are equivalences of abelian categories
${\rm Rep}(A,\xi^{-1})_{k}$$\mathscr{B}$$\mathscr{A}$$\scriptstyle{{\rm Ind}_{(A,\xi^{-1})}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}}$$\scriptstyle{{\rm Res}_{(A,\xi^{-1})}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}}$
Moreover, the functors ${\rm Ind}_{(A,\xi^{-1})}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}$ and ${\rm Res}_{(A,\xi^{-1})}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}$ are mutually inverse, and can be described explicitly. ∎
Using Theorem 3.8, we get yet another abelian equivalence, which is a generalization of [GS, Theorem 3.11], and will be used later on in §3.
Theorem 3.13.
We have an equivalence of abelian categories
${\rm F}\colon{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}\mathrm{Coh}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}(A\backslash(B\times C),\Phi),$
$V\mapsto\mathscr{O}(A\backslash(B\times C))\otimes_{k}V.$
In §4, we recall double cosets in $G$ [GS], and prove in Lemma 4.1 that Theorems 3.11, 3.13 can be applied to describe biequivariant sheaves on $(G,\omega)$ arising from double cosets (see Corollary 4.2).
In §5, we use Corollary 4.2 to describe the indecomposable exact module categories over any finite group scheme-theoretical category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$. Namely, let ${\mathscr{M}}:={\rm Coh}^{\left(\left(H,\psi\right),\left(K,\eta\right)\right)}(G,\omega)$ be an indecomposable exact module category over ${\mathscr{C}}\left(G,\omega,H,\psi\right)$. Let $Y$ be the finite scheme of $(H,K)$-double cosets in $G$ (see §4.1). For any closed point $Z\in Y(k)$, let ${\mathscr{M}}_{Z}\subset{\mathscr{M}}$ denote the full abelian subcategory consisting of all objects annihilated by the defining ideal $\mathscr{I}(Z)\subset\mathscr{O}(G)$ of $Z$. Our next result, which is a generalization of [GS, Theorem 5.3], classifies the simples of $\mathscr{M}$} using Theorem 3.11.
Theorem 5.1.
Let ${\mathscr{M}}:={\rm Coh}^{\left(\left(H,\psi\right),\left(K,\eta\right)\right)}(G,\omega)$ be as above.
  1. (1)
    For any closed point $Z\in Y(k)$ with representative closed point $g\in Z(k)$, we have an equivalence of abelian categories
    $\mathbf{Ind}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z},$
    where $L^{g}:=H\cap gKg^{-1}$ and $\xi_{g}\in Z^{2}(L^{g},\mathbb{G}_{m})$ is defined in §4.2.
  2. (2)
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$ is simple, and simple objects of $\mathscr{M}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$.
  3. (3)
    We have a direct sum decomposition of abelian categories
    ${\mathscr{M}}=\bigoplus_{Z\in Y(k)}\overline{{\mathscr{M}}_{Z}},$
    where $\overline{{\mathscr{M}}_{Z}}\subset{\mathscr{M}}$ denotes the Serre closure of ${\mathscr{M}}_{Z}$ inside $\mathscr{M}$}. ∎
Theorem 3.13 is a better tool to compute projective covers, and we have the following generalization of [GS, Theorem 5.5].
Theorem 5.3.
Let ${\mathscr{M}}:={\rm Coh}^{\left(\left(H,\psi\right),\left(K,\eta\right)\right)}(G,\omega)$ be as above.
  1. (1)
    For any closed point $Z\in Y(k)$ with representative $g\in Z(k)$, we have an equivalence of abelian categories
    $\mathbf{F}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z}.$
  2. (2)
    For any simple $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, we have
    $P_{{\mathscr{M}}}\left(\mathbf{F}_{Z}(V)\right)\cong\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V).$
As a consequence of these results, in Corollary 5.5 we obtain a classification of fiber functors on ${\mathscr{C}}\left(G,\omega,H,\psi\right)$.
In §6, we apply the results of §5 to study the abelian structure of ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$. In Theorem 6.1, we classify the simple objects of $\mathscr{C}$}, compute their Frobenius-Perron dimensions, describe how they relate under dualization, and provide an abelian decomposition for $\mathscr{C}$}. Then in Theorem 6.2, we give an alternative parametrization of the simples of $\mathscr{C}$}, use it to compute their projective covers, and deduce that $\mathscr{C}$} is unimodular if so is ${\rm Rep}_{k}(H)$. The description provided for $\mathscr{C}$} is nicer when $G$ is either étale or connected, or $H$ is normal (see §6.1–§6.3).
We conclude the paper with §7, in which we focus on the center ${\mathscr{Z}}(G,\omega)\coloneqq{\mathscr{Z}}(\rm{Coh}(G,\omega))$. Since ${\mathscr{Z}}(G,\omega)$ is a finite group scheme-theoretical category, we can apply the results from §6 to provide a description of its simple and projective objects (see Theorem 7.4). In particular, we obtain a decomposition of abelian categories
$\mathscr{Z}(G,\omega)=\bigoplus_{C\in{\rm C}(k)}\overline{\mathscr{Z}(G,\omega)_{C}},$
where $\rm C$} denotes the finite scheme of conjugacy orbits in $G$, and for each subcategory $\mathscr{Z}(G,\omega)_{C}$, we have an explicit abelian equivalence
$\mathbf{F}_{C}:{\rm Rep}_{k}(G_{C},\omega_{g})\xrightarrow{\simeq}\mathscr{Z}(G,\omega)_{C},\,\,\,V\mapsto\mathscr{O}(C)\otimes_{k}V,$
where $G_{C}$ is the stabilizer of $g\in C(k)$ and $\omega_{g}$ is defined in (2.15).
On the other hand, by [GNN, Theorem 3.5], if we set $\mathscr{D}:={\rm Coh}(G,\omega)$ and $\mathscr{D}^{\circ}:={\rm Coh}(G^{\circ},\omega^{\circ})$, there is an equivalence of tensor categories
$F:\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}=\left(\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)\right)^{G(k)}.$
This equivalence allows us to provide a more concrete description of the Serre closure $\overline{\mathscr{Z}(G,\omega)_{C}}$, generalizing [GS, Theorem 8.6].
Theorem 7.7.
For any $C\in{\rm C}(k)$, the functor $F$ restricts to an equivalence of abelian categories
$F_{C}:\overline{\mathscr{Z}(G,\omega)_{C}}\xrightarrow{\simeq}\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}.$
In particular, $F$ restricts to an equivalence of tensor categories
$F_{1}:\overline{{\rm Rep}_{k}(G)}\xrightarrow{\simeq}\mathscr{Z}\left(G^{\circ},\omega^{\circ}\right)^{G(k)}.$
2. Preliminaries
2.1. Conventions
We work over an algebraically closed field $k$ of characteristic $p\geq 0$.
All schemes $X$ considered in this paper are assumed to be finite over $k$, and equipped with a scheme morphism ${\rm Spec}(k)\to X$, unless otherwise stated. Equivalently, we will assume that $\mathscr{O}(X)$ is a finite dimensional commutative $k$-algebra with augmentation map $\varepsilon:\mathscr{O}(X)\to k$.
We assume familiarity with the theory of finite tensor categories and their modules categories, and refer to [EGNO] for any unexplained notion.
2.2. Sheaves on finite schemes
For a finite scheme $X$ as in §2.1, let ${\rm Coh}(X)$ be the abelian category of sheaves on $X$, i.e., the category of finite dimensional representations of the finite dimensional commutative algebra $\mathscr{O}(X)$. Given a closed point $x\in X(k)$, let ${\rm Coh}(X)_{x}\subset{\rm Coh}(X)$ be the abelian subcategory of sheaves on $X$ supported on $x$, so
${\rm Coh}(X)=\bigoplus_{x\in X(k)}{\rm Coh}(X)_{x}$
as abelian categories, where each ${\rm Coh}(X)_{x}$ contains a unique (up to isomorphism) simple object $\delta_{x}$ of ${\rm Coh}(X)$. We denote by $P_{x}:=P(\delta_{x})$ its projective cover, which also lies in ${\rm Coh}(X)_{x}$.
Let $Y$ be a finite scheme as in §2.1, $\varphi:Y\to X$ a scheme morphism, and $\varphi^{\sharp}:\mathscr{O}(X)\to\mathscr{O}(Y)$ the corresponding algebra homomorphism. Recall that $\varphi$ induces a pair $(\varphi^{*},\varphi_{*})$ of adjoint functors of abelian categories
(2.1) $\varphi^{*}:{\rm Coh}(X)\to{\rm Coh}(Y),\,\,\,S\mapsto S\otimes_{\mathscr{O}(X)}\mathscr{O}(Y),$
where $\mathscr{O}(X)$ acts on $\mathscr{O}(Y)$ via $\varphi^{\sharp}$, and
(2.2) $\varphi_{*}:{\rm Coh}(Y)\to{\rm Coh}(X),\,\,\,T\mapsto T_{\mid\mathscr{O}(X)},$
where $\mathscr{O}(X)$ acts on $T$ via $\varphi^{\sharp}$.
2.3. Finite group schemes
A finite group scheme $G$ is a finite scheme whose coordinate algebra $\mathscr{O}(G)$ is a (finite dimensional commutative) Hopf algebra (see, e.g., [J, W]). It is called étale if $G=G(k)$, where $G(k)$ is the group of closed points of $G$; these finite group schemes are in correspondence with finite abstract groups (see, e.g., [W, Section 6.4]). On the other hand, $G$ is connected if $G(k)=1$. For example, if $\mathfrak{g}$ is a finite dimensional $p$-Lie algebra over $k$ ($p>0$), its $p$-restricted universal enveloping algebra $u(\mathfrak{g})$ is a finite dimensional cocommutative Hopf algebra (see, e.g., [J]), and the dual commutative Hopf algebra $u(\mathfrak{g})^{*}$ is local, and satisfies $x^{p}=0$ for all $x$ in the augmentation ideal. Thus, the associated group scheme is connected. Moreover, any finite connected group scheme is obtained by successive extensions of dualized $p$-restricted enveloping algebras (see [DG, Section 2.7]).
By a theorem of Cartier (see [W, Section 11.4]), if $p=0$ then every finite group scheme is étale. On the other hand, if $p>0$ then any finite group scheme $G$ is an extension of a connected group scheme by an étale one. More precisely, $G$ fits into a split exact sequence
(2.3) $1\to G^{\circ}\xrightarrow{i}G\mathrel{\mathop{\rightleftarrows}^{\pi}_{q}}G(k)\to 1,$
where $G^{\circ}$ is connected and $G(k)$ is étale. In particular, $G(k)$ acts on ${\rm Coh}(G)_{1}\cong{\rm Coh}(G^{\circ})$, say via $a\mapsto T_{a}$, and we have an equivalence
(2.4) ${\rm Coh}(G)\simeq{\rm Coh}(G^{\circ})\rtimes G(k).$
Namely, ${\rm Coh}(G)={\rm Coh}(G^{\circ})\boxtimes{\rm Coh}(G(k))$ as abelian categories, and for any $X_{1},X_{2}\in{\rm Coh}(G^{\circ})$ and $a_{1},a_{2}\in G(k)$, we have
$\left(X_{1}\boxtimes a_{1}\right)\otimes\left(X_{2}\boxtimes a_{2}\right)=\left(X_{1}\otimes T_{a_{1}}\left(X_{2}\right)\right)\boxtimes a_{1}a_{2}.$
2.4. Quotients and free actions
Let $C$ be a finite group scheme, and $X$ a finite scheme as in §2.1, equipped with a right $C$-action
(2.5) $\mu:=\mu_{X\times C}:X\times C\to X.$
Equivalently, the algebra homomorphism
(2.6) $\mu^{\sharp}:\mathscr{O}(X)\to\mathscr{O}(X)\otimes\mathscr{O}(C)$
endows $\mathscr{O}(X)$ with a structure of a right $\mathscr{O}(C)$-comodule algebra (see [A]). Since $C$ is finite, there exists a quotient finite scheme
(2.7) $\pi:X\twoheadrightarrow X/C,$
with coordinate algebra
(2.8) $\mathscr{O}(X/C)=\mathscr{O}(X)^{C}:=\{f\in\mathscr{O}(X)\mid\mu^{\sharp}(f)=f\otimes 1\}.$
Recall that the action $\mu$ (2.5) is called free if the morphism
$(p_{1},\mu):X\times C\to X\times X$
is a closed immersion, where $p_{1}:X\times C\to X$ is the obvious projection morphism. Recall [Mu, Theorem 1(B), p.112] that in this case, the morphism $(p_{1},\mu)$ induces a scheme isomorphism
$X\times C\xrightarrow{\cong}X\times_{X/C}X.$
Equivalently, $(p_{1},\mu)$ induces an algebra isomorphism
$\mathscr{O}(X)\otimes_{\mathscr{O}(X/C)}\mathscr{O}(X)\xrightarrow{\cong}\mathscr{O}(X)\otimes_{k}\mathscr{O}(C),$
[…]
2.5. (Bi)equivariant sheaves
Retain the notation of §2.4. Let $\rm m$} be the multiplication map of $C$, and set
(2.9) $\nu:=\mu({\rm id}_{X}\times{\rm m})=\mu(\mu\times{\rm id}_{C}):X\times C\times C\to X.$
Consider the obvious projection morphisms
$p_{1}:X\times C\twoheadrightarrow X,\,\,\,q_{1}:X\times C\times C\twoheadrightarrow X,$
${\rm and}\,\,\,q_{12}:X\times C\times C\twoheadrightarrow X\times C.$
Clearly, $p_{1}\circ q_{12}=q_{1}$.
Assume we have fixed an action of $\mathscr{O}(X)$ on $\mathscr{O}(C)$ (i.e., a coaction scheme morphism $C\to X\times C$). Suppose that $\Phi\in C^{3}(X,\mathbb{G}_{m})$ and $\gamma\in C^{2}(C,\mathbb{G}_{m})$ are cochains 222 I.e., $\gamma\in\mathscr{O}(C)^{\otimes 2}$, $\Phi\in\mathscr{O}(X)^{\otimes 3}$ are invertible, $(\varepsilon\otimes{\rm id})(\gamma)=({\rm id}\otimes\varepsilon)(\gamma)=1$ and $(\varepsilon\otimes{\rm id}\otimes{\rm id})(\Phi)=({\rm id}\otimes\varepsilon\otimes{\rm id})(\Phi)=({\rm id}\otimes{\rm id}\otimes\varepsilon)(\Phi)=1\otimes 1$. such that
(2.10) $({\rm id}\otimes\Delta)(\gamma)(1\otimes\gamma)=\Phi\cdot(\Delta\otimes{\rm id})(\gamma)(\gamma\otimes 1).$
Note that multiplication by $\gamma$, and action by $\Phi$, define automorphisms of any sheaf on $C\times C$, and $X\times C\times C$, respectively, which we still denote by $\gamma$ and $\Phi$. In particular, (2.10) is equivalent to
(2.11) $d\gamma:=({\rm id}\otimes\Delta)(\gamma)(1\otimes\gamma)(\Delta\otimes{\rm id})(\gamma^{-1})(\gamma^{-1}\otimes 1)=\Phi,$
as sheaves on $X\times C\times C$.
Remark 2.1.
If $\Phi=1$, then (2.11) means that $\gamma\in Z^{2}(C,\mathbb{G}_{m})$ is a $2$-cocycle, i.e., a Drinfeld twist for $\mathscr{O}(C)$. ∎
Definition 2.2.
  1. (1)
    A right $(C,\gamma)$-equivariant sheaf on $(X,\Phi)$ is a pair $(S,\rho)$, where $S\in{\rm Coh}(X)$ and $\rho:p_{1}^{*}S\xrightarrow{\cong}\mu^{*}S$ is an isomorphism of sheaves on $X\times C$ such that the diagram
    ${q_{1}^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{({\rm id}\times{\rm m})^{*}(\rho)}$$\scriptstyle{q_{12}^{*}(\rho)}$${(\mu\circ q_{12})^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu\times{\rm id})^{*}(\rho)}$${\nu^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi\cdot({\rm id}\otimes\gamma)}$${\nu^{*}S}$
    of morphisms of sheaves on $X\times C\times C$ is commutative.
  2. (2)
    Let $(S,\rho)$, $(T,\tau)$ be two $(C,\gamma)$-equivariant sheaves on $X$. A morphism $\phi:S\to T$ in ${\rm Coh}(X)$ is $(C,\gamma)$-equivariant if the diagram
    ${p_{1}^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho}$$\scriptstyle{p_{1}^{*}(\phi)}$${p_{1}^{*}T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau}$${\mu^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu^{*}(\phi)}$${\mu^{*}T}$
    of morphisms of sheaves on $X\times C$ is commutative.
Let ${\rm Coh}^{(C,\gamma)}(X,\Phi)$ denote the category of right $(C,\gamma)$-equivariant sheaves on $(X,\Phi)$ with $(C,\gamma)$-equivariant morphisms. ∎
Remark 2.3.
Morally, a right $(C,\gamma)$-equivariant sheaf on $(X,\Phi)$ is one equipped with a lift of the $C$-action on $X$ to a $C$-action on the sheaf, which is consistent with the automorphisms of sheaves on $C\times C$ and $X\times C\times C$ induced by $\gamma$ and $\Phi$, respectively. ∎
Example 2.4.
[GS, Example 2.4] Consider $\mathscr{O}(X)\otimes\mathscr{O}(C)$ as an $\mathscr{O}(X)$-module via $\mu^{\sharp}$ (2.6). If $\gamma=1$ (so, $\Phi=1$) then $\mu^{\sharp}$ determines a $C$-equivariant structure on $\mathscr{O}(X)$. ∎
Example 2.5.
Let $\gamma\in Z^{2}(C,\mathbb{G}_{m})$, $C_{\gamma}$ the finite group scheme central extension of $C$ by $\mathbb{G}_{m}$ associated with $\gamma$, and ${\rm Rep}_{k}(C,\gamma)$ the category of finite dimensional left representations of $C_{\gamma}$ on which $\mathbb{G}_{m}$ acts trivially. Then ${\rm Rep}_{k}(C,\gamma)\cong{\rm Coh}^{(C,\gamma)}({\rm pt})$ as abelian categories. ∎
Remark 2.6.
If a finite group scheme $B$ acts on $X$ from the left, $X$ coacts on $B$, and $\beta\in C^{2}(B,\mathbb{G}_{m})$ satisfies (2.10), then the category $^{(B,\beta)}{\rm Coh}(X,\Phi)$ of left $(B,\beta)$-equivariant sheaves on $(X,\Phi)$ is defined similarly. ∎
Definition 2.7.
Suppose $X$ is equipped with a left $B$-action $\mu_{B\times X}$ and a right $C$-action $\mu_{X\times C}$, and $\Phi$, $\beta$ and $\gamma$ are as above. A $((B,\beta),(C,\gamma))$-biequivariant sheaf on $(X,\Phi)$ is a triple $(S,\lambda,\rho)$, where
$(S,\lambda)\in{}^{(B,\beta)}{\rm Coh}(X,\Phi),\,\,\,(S,\rho)\in{\rm Coh}^{(C,\gamma)}(X,\Phi),$
and the following diagram is commutative
${B\times X\times C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm id}\otimes\mu_{X\times C}}$$\scriptstyle{(\mu_{B\times X}\otimes{\rm id})\Phi}$${X\times C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{X\times C}}$${B\times X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{B\times X}}$${X.}$
The category of $((B,\beta),(C,\gamma))$-biequivariant sheaves on $(X,\Phi)$ is denoted by ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$. ∎
Note that we have
${}^{(B,\beta)}{\rm Coh}(X,\Phi)={\rm Coh}^{((B,\beta),(1,1))}(X,\Phi),\,\,\,{\rm and}$
${\rm Coh}^{(C,\gamma)}(X,\Phi)={\rm Coh}^{((1,1),(C,\gamma))}(X,\Phi).$
2.6. (Bi)equivariant morphisms
Retain the setting of §2.5. Assume that $Y$ is a finite scheme as in §2.1, on which $B$ acts from the left via $\mu_{B\times Y}$ and $C$ acts from the right via $\mu_{Y\times C}$.
Let $\varphi:Y\to X$ be a scheme morphism, and recall the adjoint functors $\varphi^{*}$ and $\varphi_{*}$ (2.1), (2.2).
Proposition 2.8.
The following hold:
  1. (1)
    If $\varphi$ is $C$-equivariant, $(\varphi^{*},\varphi_{*})$ lifts to adjoint abelian functors
    ${\rm Coh}^{(C,\gamma)}(X,\Phi)\mathrel{\mathop{\rightleftarrows}^{\varphi^{*}}_{\varphi_{*}}}{\rm Coh}^{(C,\gamma)}(Y,\varphi^{\sharp\otimes 3}\Phi).$
  2. (2)
    If $\varphi$ is $B$-equivariant, $(\varphi^{*},\varphi_{*})$ lifts to adjoint abelian functors
    $^{(B,\beta)}{\rm Coh}(X,\Phi)\mathrel{\mathop{\rightleftarrows}^{\varphi^{*}}_{\varphi_{*}}}{\rm Coh}^{(B,\beta)}(Y,\varphi^{\sharp\otimes 3}\Phi).$
  3. (3)
    If $\varphi$ is $(B,C)$-biequivariant, $(\varphi^{*},\varphi_{*})$ lifts to adjoint abelian functors
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)\mathrel{\mathop{\rightleftarrows}^{\varphi^{*}}_{\varphi_{*}}}{\rm Coh}^{((B,\beta),(C,\gamma))}(Y,\varphi^{\sharp\otimes 3}\Phi).$
Proof.
We prove (1); the proofs of (2) and (3) being similar.
Let $(S,\rho)\in{\rm Coh}^{(C,\gamma)}(X,\Phi)$, i.e., $S\in{\rm Coh}(X)$ and $\rho:S\to S\otimes\mathscr{O}(C)_{\gamma}$ endows $S$ with a structure of a right $\mathscr{O}(C)_{\gamma}$-comodule in ${\rm Coh}(X,\Phi)$. It is straightforward to verify that
$\rho^{*}:=\mu_{Y\times C}^{\sharp}\bar{\otimes}\rho:\mathscr{O}(Y)\otimes_{\mathscr{O}(X)}S\to\mathscr{O}(Y)\otimes_{\mathscr{O}(X)}S\otimes\mathscr{O}(C)_{\gamma},$
$f\otimes_{\mathscr{O}(X)}s\mapsto f^{0}\otimes_{\mathscr{O}(X)}s^{0}\otimes f^{1}s^{1},$
endows $\varphi^{*}S$ with a structure of an object $(\varphi^{*}S,\rho^{*})$ in ${\rm Coh}^{(C,\gamma)}(Y,\varphi^{\sharp\otimes 3}\Phi)$.
Let $(T,\tau)\in{\rm Coh}^{(C,\gamma)}(Y,\varphi^{\sharp\otimes 3}\Phi)$, i.e., $\tau:T\to T\otimes\mathscr{O}(C)_{\gamma}$ is a right $\mathscr{O}(C)_{\gamma}$-caction on $T$ in ${\rm Coh}^{(C,\gamma)}(Y,\varphi^{\sharp\otimes 3}\Phi)$. Then it is easy to verify that $\tau_{*}:=\tau$ endows $\varphi_{*}T$ with a structure of an object $(\varphi_{*}T,\tau_{*})$ in ${\rm Coh}^{(C,\gamma)}(X,\Phi)$.
Finally, it is straightforward to verify that the adjunction
${\rm Hom}_{{\rm Coh}(Y,\varphi^{\sharp\otimes 3}\Phi)}(\varphi^{*}S,T)\cong{\rm Hom}_{{\rm Coh}(X,\Phi)}(S,\varphi_{*}T)$
preserves $(C,\gamma)$-equivariant maps, and hence lifts to an isomorphism
${\rm Hom}_{{\rm Coh}^{(C,\gamma)}(Y,\varphi^{\sharp\otimes 3}\Phi)}((\varphi^{*}S,\rho^{*}),(T,\tau))\cong{\rm Hom}_{{\rm Coh}^{(C,\gamma)}(X,\Phi)}((S,\rho),(\varphi_{*}T,\tau_{*})),$
as desired. ∎
Example 2.9.
The scheme morphism ${\rm u}:Y\to{\rm Spec}(k)$ is $C$-equivariant, hence it induces a functor ${\rm u}^{*}:{\rm Rep}_{k}(C,\gamma)\to{\rm Coh}^{(C,\gamma)}(Y)$. ∎
2.7. The tensor category ${\rm Coh}(G,\omega)$
Let $G$ be a finite group scheme, and $\omega\in Z^{3}(G,\mathbb{G}_{m})$ a $3$-cocycle, i.e., a Drinfeld associator for $\mathscr{O}(G)$. Namely, $\omega\in\mathscr{O}(G)^{\otimes 3}$ is invertible and satisfies the equations
(2.12) $({\rm id}\otimes{\rm id}\otimes\Delta)(\omega)(\Delta\otimes{\rm id}\otimes{\rm id})(\omega)=(1\otimes\omega)({\rm id}\otimes\Delta\otimes{\rm id})(\omega)(\omega\otimes 1),$
$(\varepsilon\otimes{\rm id}\otimes{\rm id})(\omega)=({\rm id}\otimes\varepsilon\otimes{\rm id})(\omega)=({\rm id}\otimes{\rm id}\otimes\varepsilon)(\omega)=1.$
Recall [G1] that the category ${\rm Coh}(G)$ with tensor product given by convolution of sheaves and associativity constraint given by $\omega$, i.e.,
$X\otimes(Y\otimes Z)\xrightarrow{\cong}(X\otimes Y)\otimes Z,\,\,\,x\otimes y\otimes z\mapsto\omega\cdot(x\otimes y\otimes z),$
is a finite tensor category, denoted by ${\rm Coh}(G,\omega)$, and we have an equivalence of tensor categories
${\rm Coh}(G,\omega)\simeq{\rm Rep}_{k}(\mathscr{O}(G),\omega)$
with the representation category of the quasi-Hopf algebra $(\mathscr{O}(G),\omega)$.
For any closed point $g\in G(k)$, let
(2.13) $\omega_{3}^{g}:=({\rm Ad}g\otimes{\rm Ad}g)({\rm id}\otimes{\rm id}\otimes g)(\omega),$
(2.14) $\omega_{1}:=(g\otimes{\rm id}\otimes{\rm id})(\omega),\,\,\,{\rm and}\,\,\,\omega_{2}:=({\rm id}\otimes g\otimes{\rm id})(\omega),$
and set
(2.15) $\omega_{g}:=\omega_{1}\cdot({\rm Ad}g\otimes{\rm id})(\omega_{2}^{-1})\cdot\omega_{3}^{g}\in C^{2}(G,\mathbb{G}_{m}).$
For example, in the étale case, we have
$\omega_{g}(x,y)=\frac{\omega(g,x,y)\omega(gxg^{-1},gyg^{-1},g)}{\omega(gxg^{-1},g,y)};\,\,\,x,y\in G.$
2.8. (Bi)equivariant sheaves and (bi)comodules
Retain the setting of §2.5. Assume further that $X$ is a finite group scheme, and $\Phi\in\mathscr{O}(X)^{\otimes 3}$ is a Drinfeld associator, so that ${\rm Coh}(X,\Phi)$ is a finite tensor category (see §2.7).
Let $\mathscr{O}(C)_{\gamma}$ denote the vector space $\mathscr{O}(C)$ equipped with the $k$-linear map $\Delta_{\gamma}$ given by $\Delta_{\gamma}(f):=\Delta(f)\gamma$, where $\Delta$ is the standard comultiplication map of $\mathscr{O}(C)$.
Lemma 2.10.
$\mathscr{O}(C)_{\gamma}$ is a coalgebra in ${\rm Coh}(X,\Phi)$, i.e.,
$\Phi\cdot(\Delta_{\gamma}\otimes{\rm id})\Delta_{\gamma}=({\rm id}\otimes\Delta_{\gamma})\Delta_{\gamma}.$
Proof.
Follows from the compatibility (2.10). ∎
Example 2.11.
If $\gamma\in Z^{2}(C,\mathbb{G}_{m})$, then $\mathscr{O}(C)_{\gamma}$ is an ordinary coalgebra with comultiplication map $\Delta_{\gamma}$. Clearly, $\mathscr{O}(C)_{\gamma}$ is a $C$-coalgebra, which is isomorphic to the regular representation of $C$ as a $C$-module, and the category ${\rm Rep}_{k}(C,\gamma)$ (see Example 2.5) is equivalent to the category ${\rm Corep}(\mathscr{O}(C)_{\gamma})_{k}$ of finite dimensional right $k$-comodules over $\mathscr{O}(C)_{\gamma}$. ∎
Definition 2.12.
(1) Let ${\rm Comod}(\mathscr{O}(C)_{\gamma})_{{\rm Coh}(X,\Phi)}$ be the abelian category of right $\mathscr{O}(C)_{\gamma}$-comodules in ${\rm Coh}(X,\Phi)$. Explicitly, objects in this category are pairs $(S,\rho)$, where $S\in{\rm Coh}(X)$ and $\rho:S\to S\otimes\mathscr{O}(C)_{\gamma}$ is such that
$\rho(f\cdot s)=\mu^{\sharp}(f)\cdot\rho(s);\,\,\,f\in\mathscr{O}(X),\,s\in S$
(where $\mu$ is given in (2.6)), and
$(\rho\otimes{\rm id})\rho=\Phi\cdot({\rm id}\otimes\Delta_{\gamma})\rho.$
Morphisms in this category are those that preserve the actions and coactions.
(2) The abelian category ${\rm Comod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta})$ of left comodules over $\mathscr{O}(B)_{\beta}$ in ${\rm Coh}(X,\Phi)$ is defined similarly. ∎
Definition 2.13.
Let ${\rm Bicomod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta},\mathscr{O}(C)_{\gamma})$ be the abelian category of $(\mathscr{O}(B)_{\beta},\mathscr{O}(C)_{\gamma})$-bicomodules in ${\rm Coh}(X,\Phi)$. Namely, objects in this category are triples $(S,\lambda,\rho)$, where
$(S,\lambda)\in{\rm Comod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta}),\,\,(S,\rho)\in{\rm Comod}(\mathscr{O}(C)_{\gamma})_{{\rm Coh}(X,\Phi)},$
and
$(\lambda\otimes{\rm id})\rho=\Phi\cdot({\rm id}\otimes\rho)\lambda.$
Morphisms in this category are those that preserve the actions and coactions. ∎
Proposition 2.14.
The following hold:
  1. (1)
    There are $k$-linear equivalences of categories
    ${\rm Coh}^{(C,\gamma)}(X,\Phi)\simeq{\rm Comod}(\mathscr{O}(C)_{\gamma})_{{\rm Coh}(X,\Phi)},$
    ${}^{(B,\beta)}{\rm Coh}(X,\Phi)\simeq{\rm Comod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta}),\,\,\,{\rm and}\,\,\,$
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)\simeq{\rm Bicomod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta},\mathscr{O}(C)_{\gamma}).$
    In particular, the categories ${\rm Coh}^{(C,\gamma)}(X,\Phi)$, $^{(B,\beta)}{\rm Coh}(X,\Phi)$, and ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$ are abelian.
  2. (2)
    If $\mathscr{I}\subset\mathscr{O}(X)$ is a $C$-stable ideal ($B$-stable, $(B,C)$-bistable, respectively), then for any $\in{\rm Coh}^{(C,\gamma)}(X,\Phi), S\mathscr{I}S$ is a subobject of $S$ in ${\rm Coh}^{(C,\gamma)}(X,\Phi)$ ($^{(B,\beta)}{\rm Coh}(X,\Phi)$, ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$, respectively).
Proof.
(1) The proof is similar to [G1, Proposition 3.7(3)].
(2) Follows from (1) since the equivariant structure of $S$ restricts to $\mathscr{I}S$ via the $C$-equivariant ($B$-equivariant, $(B,C)$-biequivariant, respectively) inclusion morphism $\mathscr{I}S\hookrightarrow S$. ∎
2.9. Principal homogeneous spaces
Retain the notation from §2.6, §2.7. In this section we take $(G,\omega)$ for $(X,\Phi)$, assume that $\iota:H\xrightarrow{1:1}G$ is an embedding of group schemes, $\mathscr{O}(G)$ acts on $\mathscr{O}(H)$ via $\iota^{\sharp}$ and $\psi\in C^{2}(H,\mathbb{G}_{m})$ satisfies $d\psi=\iota^{\sharp\otimes 3}(\omega)$, and take $(H,\psi)$ for $(C,\gamma)$.
Consider the free right action of $H$ on $G$, given by
(2.16) $\mu_{G\times H}:G\times H\to G,\,\,\,(g,h)\mapsto gh,$
and (see §2.4) the corresponding quotient morphism
(2.17) $\pi:G\twoheadrightarrow G/H.$
Recall that $\mathscr{O}(G/H)\subset\mathscr{O}(G)$ is a right $\mathscr{O}(H)$-Hopf Galois cleft extension. Namely, $\mathscr{O}(G)$ is a right $\mathscr{O}(H)$-comodule algebra via $\mu_{G\times H}^{\sharp}$, $\mathscr{O}(G/H)\subset\mathscr{O}(G)$ is the left coideal subalgebra of coinvariants, the map
[…]
is bijective, and there exists a (unique up to multiplication by a convolution invertible element in ${\rm Hom}_{k}(\mathscr{O}(H),\mathscr{O}(H\backslash G))$) unitary convolution invertible right $\mathscr{O}(H)$-colinear map, called the cleaving map,
(2.18) $\mathfrak{c}:\mathscr{O}(H)\xrightarrow{1:1}\mathscr{O}(G).$
A choice of $\mathfrak{c}$ determines a convolution invertible $2$-cocycle
(2.19) $\sigma:\mathscr{O}(H)^{\otimes\,2}\to\mathscr{O}(G/H),\,\,\,f\otimes\tilde{f}\mapsto\mathfrak{c}(f_{1})\mathfrak{c}(\tilde{f}_{1})\mathfrak{c}^{-1}(f_{1}\tilde{f}_{2}),$
from which one can define the crossed product algebra $\mathscr{O}(G/H)\#_{\sigma}\mathscr{O}(H)$ (with the trivial action of $\mathscr{O}(H)$ on $\mathscr{O}(G/H)$), which is naturally a left $\mathscr{O}(G/H)$-module and right $\mathscr{O}(H)$-comodule algebra, where $\mathscr{O}(H)$ coacts via $a\#f\mapsto a\#f_{1}\otimes f_{2}$. Moreover, the map
(2.20) $\phi:\mathscr{O}(G/H)\#_{\sigma}\mathscr{O}(H)\xrightarrow{\cong}\mathscr{O}(G),\,\,\,a\#f\mapsto a\mathfrak{c}(f),$
is an algebra isomorphism, $\mathscr{O}(G/H)$-linear and right $\mathscr{O}(H)$-colinear. The inverse of $\phi$ is given by
(2.21) $\phi^{-1}:\mathscr{O}(G)\xrightarrow{\cong}\mathscr{O}(G/H)\#_{\sigma}\mathscr{O}(H),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\iota^{\sharp}({\rm f}_{2}))\#\iota^{\sharp}({\rm f}_{3}).$
We also have the $\mathscr{O}(G/H)$-linear map
(2.22) $\alpha:=({\rm id}\bar{\otimes}\varepsilon)\phi^{-1}:\mathscr{O}(G)\twoheadrightarrow\mathscr{O}(G/H),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\iota^{\sharp}({\rm f}_{2})).$
(For details, see [Mo, Section 3] and references therein.)
Since $\sigma$ (2.19) is a $2$-cocycle, it follows that for any $S\in{\rm Coh}(G/H)$, the vector space $S\otimes_{k}\mathscr{O}(H)$ is an $\mathscr{O}(G/H)\#_{\sigma}\mathscr{O}(H)$-module via
$(a\#f)\cdot\left(s\otimes\tilde{f}\right)=a\sigma(f_{1},\tilde{f}_{1})\cdot s\otimes f_{2}\tilde{f}_{2}.$
Consequently, we have the following lemma.
Lemma 2.15.
For any $U\in{\rm Coh}(G/H)$, the following hold:
  1. (1)
    The vector space $U\otimes_{k}\mathscr{O}(H)$ is an $\mathscr{O}(G)$-module via $\phi^{-1}$ (2.21):
    ${\rm f}\cdot\left(u\otimes\tilde{f}\right):=\phi^{-1}\left({\rm f}\right)\cdot\left(u\otimes\tilde{f}\right);\,\,{\rm f}\in\mathscr{O}(G),\,u\otimes\tilde{f}\in U\otimes\mathscr{O}(H).$
  2. (2)
    We have an $\mathscr{O}(G)$-linear isomorphism
    $F_{U}:U\otimes_{\mathscr{O}(G/H)}\mathscr{O}(G)\xrightarrow{\cong}U\otimes_{k}\mathscr{O}(H),\,\,\,u\otimes_{\mathscr{O}\left(G/H\right)}{\rm f}\mapsto\phi^{-1}\left({\rm f}\right)\cdot\left(u\otimes 1\right),$
    whose inverse given by
    $F_{U}^{-1}:U\otimes_{k}\mathscr{O}(H)\xrightarrow{\cong}U\otimes_{\mathscr{O}(G/H)}\mathscr{O}(G),\,\,\,u\otimes_{k}f\mapsto u\otimes_{\mathscr{O}(G/H)}\phi(1\otimes_{k}f).$
Proof.
Follow from the preceding remarks. ∎
For any $U\in{\rm Coh}(G/H)$, define
(2.23) $\rho^{\psi}_{U}:=(F_{U}^{-1}\otimes{\rm id})({\rm id}\otimes\Delta_{\psi})F_{U}.$
For any $(S,\rho)\in{\rm Coh}^{(H,\psi)}(G,\omega)$, define the subspace of coinvariants
(2.24) $S^{(H,\psi)}:=\{s\in S\mid\rho(s)=\mathfrak{c}(\psi^{1})\cdot s\otimes\psi^{2}\}\subset S.$
Theorem 2.16.
The following hold:
  1. (1)
    There is an equivalence of abelian categories
    ${\rm Coh}^{(H,1)}(G)\xrightarrow{\simeq}{\rm Coh}^{(H,\psi)}(G,\omega),\,\,(S,\rho)\mapsto\left(S^{(H,1)}\otimes_{k}\mathscr{O}(H),{\rm id}\otimes\Delta_{\psi}\right),$
    whose inverse is given by
    ${\rm Coh}^{(H,\psi)}(G,\omega)\xrightarrow{\simeq}{\rm Coh}^{(H,1)}(G),\,\,(S,\rho)\mapsto\left(S^{(H,\psi)}\otimes_{k}\mathscr{O}(H),{\rm id}\otimes\Delta\right)$
    (see (2.24)).
  2. (2)
    The quotient morphism $\pi:G\twoheadrightarrow G/H$ (2.17) induces an equivalence of abelian categories
    $\pi^{*}:{\rm Coh}(G/H)\xrightarrow{\simeq}{\rm Coh}^{(H,\psi)}(G,\omega),\,\,\,U\mapsto\left(U\otimes_{\mathscr{O}(G/H)}\mathscr{O}(G),\rho^{\psi}_{U}\right)$
    (see (2.23)), whose inverse is given by
    $\pi_{*}^{(H,\psi)}:{\rm Coh}^{(H,\psi)}(G,\omega)\xrightarrow{\simeq}{\rm Coh}(G/H),\,\,\,(S,\rho)\mapsto\pi^{(H,\psi)}_{*}S.$
Proof.
(1) Follows from Lemma 2.15 in a straightforward manner.
(2) For $\psi=1$, this is [Mu, p.112], so the claim follows from (1). ∎
Remark 2.17.
Note that $\omega$ plays no role in Theorem 2.16. ∎
Now for any closed point $\bar{g}:=gH\in(G/H)(k)$, let $\delta_{\bar{g}}$ denote the corresponding simple object of ${\rm Coh}(G/H)$ (see §2.2). Let
(2.25) $S_{\bar{g}}:=\pi^{*}\delta_{\bar{g}}\cong\delta_{\bar{g}}\otimes_{k}\mathscr{O}(H)\in{\rm Coh}^{(H,\psi)}(G,\omega)$
(see Theorem 2.16), and let
(2.26) $P_{\bar{g}}:=P(\delta_{\bar{g}})\in{\rm Coh}(G/H),\,\,\,P(S_{\bar{g}})\in{\rm Coh}^{(H,\psi)}(G,\omega)$
be the projective covers of $\delta_{\bar{g}}$ and $S_{\bar{g}}$, respectively.
Corollary 2.18.
The following hold:
  1. (1)
    For any $\bar{g}\in(G/H)(k)$, there are ${\rm Coh}^{(H,\psi)}(G,\omega)$-isomorphisms
    $S_{\bar{g}}\cong\delta_{g}\otimes S_{\bar{1}},\,\,\,{\rm and}\,\,\,P_{\bar{g}}\cong\delta_{g}\otimes P_{\bar{1}},$
    where $\mathscr{O}(G)$ acts diagonally and $\mathscr{O}(H)_{\psi}$ coacts on the right.
  2. (2)
    The assignment $\bar{g}\mapsto S_{\bar{g}}$ is a bijection between $(G/H)(k)$ and the set of isomorphism classes of simples in ${\rm Coh}^{(H,\psi)}(G,\omega)$.
  3. (3)
    We have $P(S_{\bar{g}})\cong\pi^{*}P_{\bar{g}}\cong P_{gH}:=\bigoplus_{h\in H(k)}P_{gh}$.
Proof.
Follow immediately from Lemma 2.15 and Theorem 2.16. ∎
Consider next the free left action of $H$ on $G$ given by
(2.27) $\mu_{H\times G}:H\times G\to G,\;\;\;(h,g)\mapsto hg,$
and the corresponding quotient morphism
(2.28) $\G.{\rm p}:G\twoheadrightarrow H\backslash$ G.
Similarly to the above, we can choose a (unique up to multiplication by an invertible element in ${\rm Hom}_{k}(\mathscr{O}(H),\mathscr{O}(H\backslash G))$) cleaving map
(2.29) $\mathfrak{c}:\mathscr{O}(H)\xrightarrow{1:1}\mathscr{O}(G),$
[…]
(2.30) $\phi:\mathscr{O}(H\backslash G)\#_{\sigma}\mathscr{O}(H)\xrightarrow{\cong}\mathscr{O}(G),\,\,\,a\#f\mapsto a\mathfrak{c}(f),$
whose inverse is given by
(2.31) $\phi^{-1}:\mathscr{O}(G)\xrightarrow{\cong}\mathscr{O}(H\backslash G)\#_{\sigma}\mathscr{O}(H),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\iota^{\sharp}({\rm f}_{2}))\#\iota^{\sharp}({\rm f}_{3}).$
[…]
(2.32) $\alpha:=({\rm id}\bar{\otimes}\varepsilon)\phi^{-1}:\mathscr{O}(G)\twoheadrightarrow\mathscr{O}(H\backslash G),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\iota^{\sharp}({\rm f}_{2})).$
It is clear that the obvious analogs of Lemma 2.15, Theorem 2.16 and Corollary 2.18 hold for $^{(H,\psi)}{\rm Coh}(G,\omega)$.
Note that the surjective Hopf algebra map $\iota^{\sharp\otimes 2}:\mathscr{O}(G)^{\otimes 2}\twoheadrightarrow\mathscr{O}(H)^{\otimes 2}$ restricts to a surjective group map $\iota^{\sharp\otimes 2}:C^{2}(G,\mathbb{G}_{m})\twoheadrightarrow C^{2}(H,\mathbb{G}_{m})$.
Lemma 2.19.
Let $\Phi\in C^{3}(G,\mathbb{G}_{m})$ be any $3$-cochain, and $\xi\in C^{2}(H,\mathbb{G}_{m})$, such that $d\xi=\iota^{\sharp\otimes 3}(\Phi)$. Assume further that $\Theta\in C^{2}(G,\mathbb{G}_{m})$ satisfies $\iota^{\sharp\otimes 2}(\Theta)=\xi$ and $\left(\iota^{\sharp\otimes 2}\otimes{\rm id}\right)(\Phi\cdot d\Theta^{-1})=1$. Then
$\left(\mathscr{O}(G),(\iota^{\sharp}\otimes{\rm id})\Delta_{\Theta}\right)\in{}^{(H,\xi)}{\rm Coh}(G,\Phi).$
Proof.
Straightforward. ∎
2.10. Twisting by the inverse $2$-cocycle
Retain the notation of §2.9. Assume that $\omega=1$, so $\psi\in Z^{2}(H,\mathbb{G}_{m})$. We note an explicit relation between the twisted coalgebras associated to $\psi$ and $\psi^{-1}$.
Set $Q_{\psi}:=\sum{\rm S}(\psi^{1})\psi^{2}$, where $\rm S$} is the antipode map of $\mathscr{O}(H)$. It is well known (see, e.g., [AEGN, Ma]) that
$Q_{\psi}^{-1}=\sum\psi^{-1}{\rm S}(\psi^{-2})\,\,\,{\rm and}\,\,\,\Delta(Q_{\psi})=({\rm S}\otimes{\rm S})(\psi_{21}^{-1})(Q_{\psi}\otimes Q_{\psi})\psi^{-1}.$
In particular, $\Delta(Q_{\psi}^{-1}{\rm S}(Q_{\psi}))=Q_{\psi}^{-1}{\rm S}(Q_{\psi})\otimes Q_{\psi}^{-1}{\rm S}(Q_{\psi})$, i.e., $Q_{\psi}^{-1}{\rm S}(Q_{\psi})$ is a grouplike element of $\mathscr{O}(H)$. Thus, we have a coalgebra isomorphism
$\mathscr{O}(H)_{\psi}^{{\rm cop}}\xrightarrow{\cong}\mathscr{O}(H)_{\psi^{-1}},\,\,\,f\mapsto{\rm S}(fQ_{\psi}^{-1}),$
and it follows that
[…]
where $\widetilde{\ell}:V\to V\otimes\mathscr{O}(H)_{\psi^{-1}}$, $v\mapsto v^{0}\otimes{\rm S}(v^{-1}Q_{\psi}^{-1})$, is an equivalence of abelian categories.
2.11. Module categories over ${\rm Coh}(G,\omega)$
Fix $(G,\omega)$ and $(H,\psi)$ as in §2.7 and §2.9, and let
(2.34) ${\mathscr{M}}(H,\psi):={\rm Coh}^{(H,\psi)}(G,\omega).$
Note that ${\mathscr{M}}(H,\psi)$ admits a canonical structure of an indecomposable exact left module category over ${\rm Coh}(G,\omega)$ given by convolution of sheaves [G1]. Namely, $X\in{\rm Coh}(G,\omega)$ acts on $(S,\rho)\in{\mathscr{M}}:={\mathscr{M}}(H,\psi)$ via $X\otimes^{{\mathscr{M}}}(S,\rho)=\left(X\otimes S,{\rm id}\otimes\rho\right)$.
Example 2.20.
The following hold:
  1. (1)
    $\mathscr{M}(1,1)\simeq{\rm Coh}(G,\omega)$ is the regular ${\rm Coh}(G,\omega)$-module.
  2. (2)
    $\mathscr{M}(G,1)\simeq{\rm Vec}$ is the usual fiber functor on ${\rm Coh}(G)$.
For any closed point $g\in G(k)$, set
(2.35) $\psi_{g}:=(\psi^{g}\omega_{g})_{\mid g^{-1}Hg}\in C^{2}(g^{-1}Hg,\mathbb{G}_{m}),$
where $\omega_{g}$ is defined in (2.15).
Definition 2.21.
Two pairs $(H,\psi)$, $(H^{\prime},\psi^{\prime})$ as above are equivalent if there is a closed point $g\in G(k)$ such that $H^{\prime}=g^{-1}Hg$ and the class of $(\psi^{\prime})^{-1}\psi_{g}$ in $H^{2}(H^{\prime},\mathbb{G}_{m})$ is trivial. ∎
Theorem 2.22.
[G1] The following hold:
  1. (1)
    There is a bijection between equivalence classes of pairs $(H,\psi)$ (in the sense of Definition 2.21) and equivalence classes of indecomposable exact left module categories over ${\rm Coh}(G,\omega)$, assigning $(H,\psi)$ to $\mathscr{M}(H,\psi)$.
  2. (2)
    The abelian equivalence $\mathscr{M}(H,\psi)\simeq{\rm Comod}(\mathscr{O}(H)_{\psi})_{{\rm Coh}(G,\omega)}$ given in Proposition 2.14 is a ${\rm Coh}(G,\omega)$-equivalence.
Remark 2.23.
If $G$ is connected, i.e., $G(k)=1$, Theorem 2.22 implies that equivalence classes of indecomposable exact module categories over ${\rm Coh}(G,\omega)$ correspond bijectively to pairs $(H,\psi)$. ∎
Remark 2.24.
Let ${\mathscr{M}}:={\mathscr{M}}(H,\psi)$. Corollary 2.18(1) states that
$S_{\bar{g}}\cong\delta_{g}\otimes^{{\mathscr{M}}}S_{\bar{1}}\,\,\,{\rm and}\,\,\,P_{\bar{g}}\cong\delta_{g}\otimes^{{\mathscr{M}}}P_{\bar{1}}$
for any simple $S_{\bar{g}}\in{\mathscr{M}}$. Also, using Corollary 2.18(3), we see that
$\begin{split}&P_{1}\otimes^{{\mathscr{M}}}S_{\bar{1}}\cong P_{1}\otimes^{{\mathscr{M}}}\mathscr{O}(H)\cong\pi_{*}^{(H,\psi)}\pi^{*}\left(P_{1}\otimes\mathscr{O}(H)\right)\\ &\cong\pi_{*}^{(H,\psi)}P_{H}\cong|H^{\circ}|P_{\bar{1}},\end{split}$
which in particular demonstrates that $\mathscr{M}$} is exact (see [G1]). ∎
Remark 2.25.
For any pair $(H,\psi)$ as above, let
$\mathscr{N}(H,\psi):={}^{(H,\psi)}{\rm Coh}(G,\omega)\simeq{\rm Comod}_{{\rm Coh}(G,\omega)}(\mathscr{O}(H)_{\psi})$
(see Proposition 2.14). The assignment $(H,\psi)\mapsto\mathscr{N}(H,\psi)$ classifies exact indecomposable right module categories over ${\rm Coh}(G,\omega)$, up to equivalence. ∎
2.12. Module categories over ${\mathscr{C}}(G,\omega,H,\psi)$
Fix $G,\omega,H,\psi$ as in §2.11. Recall [G1] that the group scheme-theoretical category
${\mathscr{C}}(G,\omega,H,\psi):={\rm Coh}(G,\omega)^{*}_{\mathscr{M}(H,\psi)}$
associated to this data is the dual category of ${\rm Coh}(G,\omega)$ with respect to $\mathscr{M}(H,\psi)$. In other words,
(2.36) ${\mathscr{C}}(G,\omega,H,\psi)={\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(H,\psi)\right)$
is the category of ${\rm Coh}(G,\omega)$-module endofunctors of $\mathscr{M}(H,\psi)$. Recall that ${\mathscr{C}}(G,\omega,H,\psi)$ is a finite tensor category with tensor product given by composition of module functors [EO].
Example 2.26.
The following hold:
${\mathscr{C}}\left(G,\omega,1,1\right)\simeq{\rm Coh}(G,\omega)\,\,\,{\rm and}\,\,\,{\mathscr{C}}\left(G,1,G,1\right)\simeq{\rm Rep}_{k}(G).$
Theorem 2.27.
[G1] Fix ${\mathscr{C}}:={\mathscr{C}}(G,\omega,H,\psi)$. The assignment
$\mathscr{M}(K,\eta)\mapsto{\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)$
determines an equivalence between the $2$-category of indecomposable exact left module categories over ${\rm Coh}(G,\omega)$ and the $2$-category of indecomposable left exact $\mathscr{C}$}-module categories. ∎
For $(K,\eta)$ as above, consider the abelian category
(2.37) $\mathscr{M}\left((H,\psi),(K,\eta)\right):={\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(G,\omega)$
of biequivariant sheaves on $(G,\omega)$ with respect to the actions $\mu_{H\times G}$ and $\mu_{G\times K}$ (see Definition 2.2).
Lemma 2.28.
The following hold:
  1. (1)
    There is an equivalences of abelian categories
    $\mathscr{M}\left((H,\psi),(K,\eta)\right)\simeq{\rm Bicomod}_{{\rm Coh}(G,\omega)}\left(\mathscr{O}(H)_{\psi},\mathscr{O}(K)_{\eta}\right).$
  2. (2)
    There is an equivalence of abelian categories
    ${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\simeq\mathscr{M}\left((H,\psi),(K,\eta)\right).$
Proof.
(1) Follows from Proposition 2.14.
(2) A functor $\mathscr{M}(H,\psi)\to\mathscr{M}(K,\eta)$ is determined by a $(K,\eta)$-equivariant sheaf $S$ on $(G,\omega)$ (its value on $\mathscr{O}(H)_{\psi}$), and the fact that the functor is a ${\rm Coh}(G,\omega)$-module (hence, ${\rm Coh}(H)$-module) functor gives $S$ a commuting left $H$-equivariant structure for the left action $\mu_{H\times G}$ of $H$ on $G$, i.e., $S\in\mathscr{M}\left((H,\psi),(K,\eta)\right)$.
Conversely, it is clear that any $S$ in $\mathscr{M}\left((H,\psi),(K,\eta)\right)$, viewed as an object in ${\rm Bicomod}_{{\rm Coh}(G,\omega)}\left(\mathscr{O}(H)_{\psi},\mathscr{O}(K)_{\eta}\right)$, defines a ${\rm Coh}(G,\omega)$-module functor
$\mathscr{M}(H,\psi)\to\mathscr{M}(K,\eta),\,\,\,T\mapsto T\otimes^{\mathscr{O}(H)_{\psi}}S,$
where $T\otimes^{\mathscr{O}(H)_{\psi}}S$ is the cotensor product. ∎
Recall that convolution of sheaves on $G$ lifts to endow the category
${\mathscr{M}}\left((H,\psi),(H,\psi)\right)={\rm Coh}^{((H,\psi),(H,\psi))}(G,\omega)$
with the structure of a finite tensor category [G1].
Lemma 2.29.
There are equivalences of tensor categories
${\mathscr{C}}(G,\omega,H,\psi)\simeq{\rm Bicomod}_{{\rm Coh}(G,\omega)}(\mathscr{O}(H)_{\psi})^{{\rm rev}},$
${\mathscr{C}}(G,\omega,H,\psi)\simeq{\mathscr{M}}\left((H,\psi),(H,\psi)\right)^{{\rm rev}}.$
Proof.
For the first equivalence see, e.g., [EGNO]. The second equivalence follows from Lemma 2.28. ∎
3. Associated biequivariant sheaves
Let $B$ and $C$ be finite group schemes. Let
(3.1) $j_{1}:B\xrightarrow{b\mapsto(b,1)}B\times C\,\,\,{\rm and}\,\,\,j_{2}:C\xrightarrow{c\mapsto(1,c)}B\times C$
be the canonical group scheme embeddings. Suppose that
(3.2) $\partial_{1}:A\xrightarrow{1:1}B\,\,\,{\rm and}\,\,\,\partial_{2}:A\xrightarrow{1:1}C$
are two group scheme embeddings, and let
(3.3) $\partial:=\left(\partial_{1},\partial_{2}\right):A\xrightarrow{1:1}B\times C,\,\,\,a\mapsto(\partial_{1}(a),\partial_{2}(a)).$
Let $A\backslash(B\times C)$ be the quotient scheme with respect to $\mu_{A\times(B\times C)}$ (2.27), and let
(3.4) ${\rm p}:B\times C\twoheadrightarrow A\backslash(B\times C)$
be the quotient morphism (see 2.28).
Fix a $3$-cochain
(3.5) $\Phi=\sum\Phi^{1}\otimes\Phi^{2}\otimes\Phi^{3}\in C^{3}(A\backslash(B\times C),\mathbb{G}_{m}).$
Suppose that $\beta\in C^{2}(B,\mathbb{G}_{m})$ and $\gamma\in C^{2}(C,\mathbb{G}_{m})$ satisfy
$d\beta=({\rm p}j_{1})^{\sharp\otimes 3}(\Phi)\,\,\,{\rm and}\,\,\,d\gamma=({\rm p}j_{2})^{\sharp\otimes 3}(\Phi),$
and set
(3.6) $\Psi:=\beta^{-1}\times\gamma\in C^{2}(B\times C,\mathbb{G}_{m}).$
Assume for the rest of this section that there exists a $2$-cochain
${\rm W}\in C^{2}(B\times C,\mathbb{G}_{m})$
such that, setting $\Theta:=\Psi^{-1}{\rm W}$, we have
(3.7) $\xi:=\partial^{\sharp\otimes 2}\left(\Theta\right)\in Z^{2}(A,\mathbb{G}_{m}).$
Example 3.1.
If $\Phi=1$, then we can take ${\rm W}=1$. In this case, we have $\Theta=\Psi^{-1}\in Z^{2}(B\times C,\mathbb{G}_{m})$, so $\xi=\partial^{\sharp\otimes 2}\left(\Theta\right)\in Z^{2}(A,\mathbb{G}_{m})$ is automatic. ∎
Consider now the right action of $C$ on $A\backslash(B\times C)$ given by
(3.8) $A\backslash(B\times C)\times C\to A\backslash(B\times C),\,\,\,\overline{(b,c_{1})}\cdot c_{2}=\overline{(b,c_{1}c_{2})},$
and left action of $B$ on $A\backslash(B\times C)$ given by
(3.9) $B\times A\backslash(B\times C)\to A\backslash(B\times C),\,\,\,b_{1}\cdot\overline{(b_{2},c)}=\overline{(b_{2}b_{1}^{-1},c)}.$
Consider also the actions of $\mathscr{O}(A\backslash(B\times C))$ on $\mathscr{O}(B)$ and $\mathscr{O}(C)$ via $({\rm p}j_{1})^{\sharp}$ and $({\rm p}j_{2})^{\sharp}$, respectively (see (3.1), (3.4)). Finally, let
(3.10) $\mathscr{A}:={\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)$
be the associated abelian category (see Definition 2.7).
Next consider the right action of $C$ on $B\times C$ given by
(3.11) $(B\times C)\times C\to B\times C,\,\,\,(b,c_{1})\cdot c_{2}=(b,c_{1}c_{2}),$
and left action of $A\times B$ on $B\times C$ given by
(3.12) $A\times B\times(B\times C)\to B\times C,\,\,\,(a,b_{1})\cdot(b_{2},c)=\partial(a)(b_{2}b_{1}^{-1},c).$
Consider also the action of $\mathscr{O}(B\times C)$ on $\mathscr{O}(A\times B)$ via the Hopf algebra map
(3.13) $\varphi:=(\partial^{\sharp}\otimes j_{1}^{\sharp})\Delta_{\mathscr{O}(B\times C)}:\mathscr{O}(B\times C)\to\mathscr{O}(A\times B),$
and the action of $\mathscr{O}(B\times C)$ on $\mathscr{O}(C)$ via $j_{2}^{\sharp}$. Finally, let
(3.14) $\mathscr{B}:={\rm Coh}^{((A\times B,\xi^{-1}\times\beta),(C,\gamma))}\left(B\times C,\Phi\right)$
be the associated abelian category (see Definition 2.7).
Let ${\rm Rep}(A,\xi^{-1})_{k}={\rm Corep}_{k}(\mathscr{O}(A)_{\xi^{-1}})$ be the category of finite dimensional left $\mathscr{O}(A)_{\xi^{-1}}$-comodules. Our goal in this section is to construct explicit equivalences of abelian categories
${\rm Rep}(A,\xi^{-1})_{k}$$\mathscr{B}$$\hskip 1.42271pt\mathscr{A}.$$\scriptstyle{{\rm t}^{*}}$$\scriptstyle{{\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\eta))}}$$\scriptstyle{{\rm t}_{*}^{((B,\beta),(C,\gamma))}}$$\scriptstyle{{}^{(A,\xi^{-1})}{\rm p}_{*}}$$\scriptstyle{{\rm p}^{*}}$$\scriptstyle{{\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}}$
3.1. The equivalence $\mathscr{B}\simeq{\rm Rep}(A,\xi^{-1})_{k}$
For any $(S,\lambda,\rho)\in\mathscr{B}$ (3.14), define the maps
(3.15) $\lambda_{1}:=(\varepsilon\otimes{\rm id}\otimes{\rm id})\lambda:S\to\mathscr{O}(B)_{\beta}\otimes S,$
(3.16) $\lambda_{2}:=({\rm id}\otimes\varepsilon\otimes{\rm id})\lambda:S\to\mathscr{O}(A)_{\xi^{-1}}\otimes S,$
(3.17) $\lambda_{\beta}:=(\beta\otimes 1)\cdot(j_{1}^{\sharp}\otimes{\rm id})\Delta_{\mathscr{O}(B\times C)},\,\,\,{\rm and}\,\,\,\rho_{\gamma}:=(1\otimes\gamma)\cdot({\rm id}\otimes j_{2}^{\sharp})\Delta_{\mathscr{O}(B\times C)},$
and the bicoinvariants subsheaf
$\begin{split}&S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)}:=\\ &\{s\in S\mid\lambda_{1}(s)=\beta^{1}\otimes(\beta^{2}\otimes 1)\cdot s,\,\,\rho(s)=(1\otimes\gamma^{1})\cdot s\otimes\gamma^{2}\}\subset S.\end{split}$
Lemma 3.2.
For any $(S,\lambda,\rho)\in\mathscr{B}$, the following hold:
  1. (1)
    $(S,\lambda_{1},\rho)\in{\rm Coh}^{((B,\beta),(C,\gamma))}(B\times C,\Phi)$.
  2. (2)
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(B\times C,\Phi)\simeq{\rm Vec}$ with the unique simple object being $\left(\mathscr{O}(B\times C),\lambda_{\beta},\rho_{\gamma}\right)$.
  3. (3)
    There is a ${\rm Coh}^{((B,\beta),(C,\gamma))}(B\times C,\Phi)$-isomorphism
    $(S,\lambda_{1},\rho)\cong\left(\mathscr{O}(B\times C),\lambda_{\beta},\rho_{\gamma}\right)\otimes_{k}S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)}.$
  4. (4)
    $(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2})\subset(S,\lambda_{2})$ in ${\rm Rep}(A,\xi^{-1})_{k}$.
Proof.
(1) is clear, (2) follows from [G1], and (3)-(4) from (1)-(2). ∎
Consider the trivial morphism ${\rm t}:B\times C\twoheadrightarrow 1$.
Theorem 3.3.
The functor ${\rm t}^{*}:{\rm Vec}\to{\rm Coh}(B\times C)$ (see §2.2) lifts to an equivalence of abelian categories
${\rm t}^{*}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}\mathscr{B},\,\,\,(V,\ell)\mapsto\left(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right)$
(see (3.17)-(3.18) below), whose inverse functor is given by
${\rm t}_{*}^{((B,\beta),(C,\gamma),\lambda_{1},\rho)}:\mathscr{B}\xrightarrow{\simeq}{\rm Rep}(A,\xi^{-1})_{k},\,\,\,(S,\lambda,\rho)\mapsto\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right).$
Proof.
Let $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, and write $\ell(v)=\sum v^{-1}\otimes v^{0}$. Consider the free $\mathscr{O}(B\times C)$-module ${\rm t}^{*}V=\mathscr{O}(B\times C)\otimes_{k}V$, and map
$\lambda^{(\ell,\beta)}:=\left(\left(1\otimes\beta\otimes 1\otimes 1\right)(\varphi\otimes{\rm id})\Delta_{\mathscr{O}(B\times C)}\right)\bar{\otimes}\ell.$
We have
(3.18) $\begin{split}&\lambda^{(\ell,\beta)}:\mathscr{O}(B\times C)\otimes_{k}V\to\mathscr{O}(A\times B)_{\xi^{-1}\times\beta}\otimes\mathscr{O}(B\times C)\otimes_{k}V,\\ &{\rm f}\otimes v\mapsto\partial^{\sharp}({\rm f}_{1})v^{-1}\otimes\beta^{1}j_{1}^{\sharp}({\rm f}_{2})\otimes(\beta^{2}\otimes 1){\rm f}_{3}\otimes v^{0}.\end{split}$
It is straightforward to verify that
$(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V})\in{\rm Coh}^{((A\times B,\xi^{-1}\times\beta),(C,\gamma))}\left(B\times C,\Phi\right).$
Thus, we have a functor
${\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{}\mathscr{B},\,\,\,(V,\ell)\mapsto\left(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right).$
Conversely, take any $(S,\lambda,\rho)\in\mathscr{B}$, and consider the sheaf ${\rm t}_{*}S$ (that is, the underlying vector space of $S$). Then by Lemma 3.2(4), we have $\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right)\in{\rm Rep}(A,\xi^{-1})_{k}$. Thus, we have a functor
$\mathscr{B}\xrightarrow{}{\rm Rep}(A,\xi^{-1})_{k},\,\,\,(S,\lambda,\rho)\mapsto\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right).$
Finally, it is straightforward to verify that the two functors constructed above are inverse to each other. ∎
Remark 3.4.
If $A$, $B$ and $C$ are any affine group schemes, then Theorem 3.3 and its proof hold after replacing $\rm Coh$} by ${\rm Coh}_{{\rm f}}$ (see [G1, Definition 3.2]). ∎
3.2. The equivalence $\mathscr{A}\simeq\mathscr{B}$
Recall that we assume that
(3.19) $\Theta:=\Psi^{-1}{\rm W}\in C^{2}(B\times C,\mathbb{G}_{m})$
satisfies (3.7). For $(S,\lambda,\rho)\in\mathscr{B}$ (3.14), define the coinvariants subsheaf
$S^{((A,\xi^{-1}),\lambda_{2})}:=\{s\in S\mid\lambda_{2}(s)=\partial^{\sharp}(\Theta^{-1})\otimes\Theta^{-2}\cdot s\}\subset S,$
where $\lambda_{2}$ is given in (3.16).
Lemma 3.5.
For any $(S,\lambda,\rho)\in\mathscr{B}$, the following hold:
  1. (1)
    $(S,\lambda_{2},\rho)\in{\rm Coh}^{((A,\xi^{-1}),(C,\gamma))}(B\times C,\Phi)$.
  2. (2)
    $\left({\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}S,\lambda_{1},\rho\right)\in\mathscr{A}$ (3.10).
Proof.
Similar to the proof of Lemma 3.2. ∎
Theorem 3.6.
The functor ${\rm p}^{*}:{\rm Coh}(A\backslash(B\times C))\to{\rm Coh}(B\times C)$ (see §2.2 and (3.4)) lifts to an equivalence of abelian categories
${\rm p}^{*}:\mathscr{A}\xrightarrow{\simeq}\mathscr{B},\,\,\,(S,\lambda,\rho)\mapsto\left(\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C))}S,\lambda^{*},\rho^{*}\right)$
(see (3.20)-(3.21) below), whose inverse is given by
${\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}:\mathscr{B}\xrightarrow{\simeq}\mathscr{A},\,\,\,(S,\lambda,\rho)\mapsto({\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}S,\lambda_{1},\rho).$
Proof.
Let $(S,\lambda,\rho)\in\mathscr{A}$ (3.10). Write
$\lambda(s)=\sum s^{-1}\otimes s^{0}\in\mathscr{O}(B)_{\beta}\otimes S,\,\,\,{\rm and}\,\,\,\rho(s)=\sum s^{0}\otimes s^{1}\in S\otimes\mathscr{O}(C)_{\gamma}.$
Then ${\rm p}^{*}S=\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C))}S$ acquires a natural structure of an object in $\mathscr{B}$ (3.14), given by
(3.20) $\begin{split}&\lambda^{*}:{\rm p}^{*}S\to\mathscr{O}(A\times B)_{\xi^{-1}\times\beta}\otimes{\rm p}^{*}S,\\ &{\rm f}\otimes s\mapsto\partial^{\sharp}(\Theta^{1}{\rm f}_{1})\otimes j_{1}^{\sharp}({\rm f}_{2})s^{-1}\otimes\Theta^{2}{\rm f}_{3}\otimes s^{0}\end{split}$
(i.e., $\lambda^{*}:=\left(\left(\partial^{\sharp}(\Theta^{1})\otimes\Theta^{2}\right)(\varphi\otimes{\rm id})\Delta_{\mathscr{O}(B\times C)}\right)\bar{\otimes}\lambda$ (3.13), (3.19)), and
(3.21) $\rho^{*}:{\rm p}^{*}S\to{\rm p}^{*}S\otimes\mathscr{O}(C)_{\gamma},\,\,\,{\rm f}\otimes s\mapsto{\rm f}_{1}\otimes s^{0}\otimes j_{2}^{\sharp}({\rm f}_{2})s^{1}.$
Thus, we have a functor
$\mathscr{A}\xrightarrow{}\mathscr{B},\,\,\,(S,\lambda,\rho)\mapsto(\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C))}S,\lambda^{*},\rho^{*}).$
Conversely, take any $(S,\lambda,\rho)\in\mathscr{B}$. Then by Lemma 3.5, we have $\left({\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}S,\lambda_{1},\rho\right)\in\mathscr{A}$, so we have a functor
$\mathscr{B}\xrightarrow{}\mathscr{A},\,\,\,(S,\lambda,\rho)\mapsto({\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}S,\lambda_{1},\rho).$
Finally, it is straightforward to verify that the two functors constructed above are inverse to each other. ∎
Remark 3.7.
If $A$, $B$ and $C$ are affine group schemes such that the quotient $A\backslash(B\times C)$ is affine, then Theorem 3.6 and its proof hold after replacing $\rm Coh$} by ${\rm Coh}_{{\rm f}}$ (see [G1, Definition 3.2]).
3.3. The first equivalence ${\rm Rep}(A,\xi^{-1})_{k}\simeq\mathscr{A}$
Set
(3.22) ${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:={\rm p}{}^{(A,\xi^{-1})}_{*}{\rm t}^{*}\,\,\,{\rm and}\,\,\,{\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:=\rm t_{*}^{((B,\beta),(C,\gamma))}{\rm p}^{*}.$
Theorem 3.8.
The equivalences
${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
$(V,\ell)\mapsto\left({\rm p}_{*}^{(A,\xi^{-1},\lambda^{(\ell,\beta)}_{2})}\left(\mathscr{O}(B\times C)\otimes_{k}V\right),\lambda^{(\ell,\beta)}_{1},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right),$
and
${\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)\xrightarrow{\simeq}{\rm Rep}(A,\xi^{-1})_{k},$
$(S,\lambda,\rho)\mapsto\left(\left(\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C))}S\right)^{((B,\beta),(C,\gamma),\lambda^{*}_{1},\rho^{*})},\lambda^{*}_{2}\right)$
(see (3.20)-(3.21)), are inverse to each other.
Proof.
Follows from Theorems 3.3 and 3.6. ∎
Now recall (3.19), and consider the linear map
(3.23) $\lambda:\mathscr{O}(B\times C)\to\mathscr{O}(A)_{\xi}\otimes\mathscr{O}(B\times C),\,\,\,{\rm f}\mapsto\partial^{\sharp}\left(\Theta^{1}{\rm f}_{1}\right)\otimes\Theta^{2}{\rm f}_{2}.$
Remark 3.9.
If $\Phi=1$ then $\Theta=\Psi^{-1}\in Z^{2}(B\times C,\mathbb{G}_{m})$ (see Example 3.1), so by Lemma 2.19, $\lambda$ is a coaction. However, if $\Phi\neq 1$ then $\lambda$ may not be a coaction (see Lemma 4.1(4) below). ∎
Set $Q:=Q_{\xi}$, so $Q={\rm S}\left(Q_{\xi^{-1}}^{-1}\right)$ (2.10). Recall (2.33) that for any $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, we have $(V,\widetilde{\ell})\in{\rm Rep}_{k}(A,\xi)$, where
$\widetilde{\ell}:V\to V\otimes\mathscr{O}(A)_{\xi},\,\,\,v\mapsto v^{0}\otimes{\rm S}\left(v^{-1}\right)Q.$
For any $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, consider the subspace
(3.24) $V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C):={\rm Ker}\left(\widetilde{\ell}\otimes{\rm id}-{\rm id}\otimes\lambda\right).$
Proposition 3.10.
For any $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, we have
$V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C)=\left(\mathscr{O}(B\times C)\otimes_{k}V\right)^{(A,\xi^{-1},\lambda_{2}^{(\ell,\beta)})}.$
Proof.
By Theorem 3.3, $\mathscr{O}(A)_{\xi^{-1}}$ coacts on $V\otimes_{k}\mathscr{O}(B\times C)$ via
$\lambda_{2}^{(\ell,\beta)}\left(v\otimes{\rm f}\right)=v^{-1}\partial^{\sharp}({\rm f}_{1})\otimes v^{0}\otimes{\rm f}_{2},$
so by definition, $\sum_{i}v_{i}\otimes{\rm f}_{i}\in\left(V\otimes_{k}\mathscr{O}(B\times C)\right)^{(A,\xi^{-1},\lambda_{2}^{(\ell,\beta)})}$ if and only if
(3.25) $\sum_{i}v_{i}^{-1}\partial^{\sharp}({\rm f}_{i1})\otimes v_{i}^{0}\otimes{\rm f}_{i2}=\sum_{i}\partial^{\sharp}(\Theta^{-1})\otimes v_{i}\otimes\Theta^{-2}{\rm f}_{i}.$
Now assume that $\sum_{i}v_{i}\otimes{\rm f}_{i}\in\left(V\otimes_{k}\mathscr{O}(B\times C)\right)^{(A,\xi^{-1},\lambda_{2}^{(\ell,\beta)})}$. We have to verify that $\sum_{i}v_{i}\otimes{\rm f}_{i}$ lies in $V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C)$, i.e., that
$\sum_{i}\widetilde{\ell}(v_{i})\otimes{\rm f}_{i}=\sum_{i}v_{i}\otimes\lambda({\rm f}_{i}).$
Thus, we have to show that
(3.26) $\sum_{i}v_{i}^{0}\otimes{\rm S}(v_{i}^{-1})Q\otimes{\rm f}_{i}=\sum_{i}v_{i}\otimes\partial^{\sharp}\left(\Theta^{1}{\rm f}_{i1}\right)\otimes\Theta^{2}{\rm f}_{i2}.$
To this end, apply ${\rm id}\otimes\widetilde{\ell}\otimes{\rm id}$ to (3.25) to obtain
$\sum_{i}v_{i}^{-2}\partial^{\sharp}({\rm f}_{i1})\otimes v_{i}^{0}\otimes{\rm S}(v_{i}^{-1})Q\otimes{\rm f}_{i2}=\sum_{i}\partial^{\sharp}(\Theta^{-1})\otimes v_{i}^{0}\otimes{\rm S}(v_{i}^{-1})Q\otimes\Theta^{-2}{\rm f}_{i}.$
Multiplying the first factor by the third one, yields
$\sum_{i}{\rm S}\left(v_{i}^{-1}\right)v_{i}^{-2}\partial^{\sharp}({\rm f}_{i1})\otimes v_{i}^{0}\otimes{\rm f}_{i2}=\sum_{i}{\rm S}\left(v_{i}^{-1}\right)\partial^{\sharp}(\Theta^{-1})\otimes v_{i}^{0}\otimes\Theta^{-2}{\rm f}_{i},$
or equivalently, since ${\rm S}\left(v_{i}^{-1}\right)v_{i}^{-2}=\varepsilon\left(v_{i}^{-1}\right)\xi^{-1}{\rm S}(\xi^{-2})=\varepsilon\left(v_{i}^{-1}\right)Q^{-1}$,
$\sum_{i}\varepsilon\left(v_{i}^{-1}\right)Q^{-1}\partial^{\sharp}({\rm f}_{i1})\otimes v_{i}^{0}\otimes{\rm f}_{i2}=\sum_{i}{\rm S}\left(v_{i}^{-1}\right)\partial^{\sharp}(\Theta^{-1})\otimes v_{i}^{0}\otimes\Theta^{-2}{\rm f}_{i}.$
Thus, we have
$\sum_{i}Q^{-1}\partial^{\sharp}({\rm f}_{i1})\otimes v_{i}\otimes{\rm f}_{i2}=\sum_{i}{\rm S}\left(v_{i}^{-1}\right)\partial^{\sharp}(\Theta^{-1})\otimes v_{i}^{0}\otimes\Theta^{-2}{\rm f}_{i},$
which is equivalent to (3.26).
Similarly, if $\sum_{i}v_{i}\otimes{\rm f}_{i}$ is in $V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C)$, then $\sum_{i}v_{i}\otimes{\rm f}_{i}$ lies in $\left(\mathscr{O}(B\times C)\otimes_{k}V\right)^{(A,\xi^{-1},\lambda_{2}^{(\ell,\beta)})}$. ∎
Theorem 3.11.
The equivalences
${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
$(V,\ell)\mapsto\left(V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C),\lambda^{(\ell,\beta)}_{1},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right),$
and
${\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)\xrightarrow{\simeq}{\rm Rep}(A,\xi^{-1})_{k},$
$(S,\lambda,\rho)\mapsto\left(\left(\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C)))}S\right)^{((B,\beta),(C,\gamma),\lambda^{*}_{1},\rho^{*})},\lambda^{*}_{2}\right),$
are inverse to each other.
Proof.
Follows from Theorem 3.8 and Proposition 3.10. ∎
Example 3.12.
Recall (3.17). There is a canonical isomorphism
${\rm Ind}\left(\mathscr{O}(A),\Delta_{\xi^{-1}}\right)\cong\left(\mathscr{O}(B\times C),\lambda_{\beta},\rho_{\gamma}\right)$
in ${\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)$. ∎
3.4. The second equivalence ${\rm Rep}(A,\xi^{-1})_{k}\simeq\mathscr{A}$
Choose a cleaving map (2.29) $\mathfrak{c}:\mathscr{O}(A)\xrightarrow{1:1}\mathscr{O}(B\times C)$. Recall (2.32), and let
$\alpha:\mathscr{O}(B\times C)\twoheadrightarrow\mathscr{O}(A\backslash(B\times C)),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\partial^{\sharp}({\rm f}_{2})).$
Also, for any $V\in{\rm Rep}(A,\xi^{-1})_{k}$, consider the $k$-linear isomorphisms
${\rm F}_{V}:=\left({\rm id}\otimes\alpha\right)_{21}:V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C)\xrightarrow{\cong}\mathscr{O}(A\backslash(B\times C))\otimes_{k}V,$
and
${\rm F}_{V}^{-1}:\mathscr{O}(A\backslash(B\times C))\otimes_{k}V\xrightarrow{\cong}V\otimes^{\mathscr{O}(A)_{\xi}}\mathscr{O}(B\times C),$
${\rm f}\otimes v\mapsto v^{0}\otimes{\rm f}\mathfrak{c}(v^{-1}).$
Recall the maps $\lambda_{\beta}$ and $\rho_{\gamma}$ (3.17), and define the maps
$\lambda_{V}:=({\rm id}\otimes{\rm F}_{V})(12)({\rm id}\otimes\lambda_{\beta}){\rm F}_{V}^{-1},\,\,\,{\rm and}\,\,\,\rho_{V}:=({\rm F}_{V}\otimes{\rm id})({\rm id}\otimes\rho_{\gamma}){\rm F}_{V}^{-1}.$
Note that we have
(3.27) $\begin{split}&\lambda_{V}:\mathscr{O}(A\backslash(B\times C))\otimes_{k}V\to\mathscr{O}(B)_{\beta}\otimes\mathscr{O}(A\backslash(B\times C))\otimes_{k}V,\\ &{\rm f}\otimes v\mapsto\beta^{1}j_{1}^{\sharp}\left({\rm f}_{1}\mathfrak{c}\left(v^{-1}\right)_{1}\right)\otimes v^{0}\otimes\alpha\left(\left(\beta^{2}\otimes 1\right){\rm f}_{2}\mathfrak{c}\left(v^{-1}\right)_{2}\right),\end{split}$
and
(3.28) $\begin{split}&\rho_{V}:\mathscr{O}(A\backslash(B\times C))\otimes_{k}V\to\mathscr{O}(A\backslash(B\times C))\otimes_{k}V\otimes\mathscr{O}(C)_{\gamma},\\ &{\rm f}\otimes v\mapsto\alpha\left(\left(1\otimes\gamma^{1}\right){\rm f}_{1}\mathfrak{c}\left(v^{-1}\right)_{1}\right)\otimes v^{0}\otimes\gamma^{2}j_{2}^{\sharp}\left({\rm f}_{2}\mathfrak{c}\left(v^{-1}\right)_{2}\right).\end{split}$
Theorem 3.13.
We have an equivalence of abelian categories
${\rm F}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
$V\mapsto(\mathscr{O}(A\backslash(B\times C))\otimes_{k}V,\lambda_{V},\rho_{V}).$
Proof.
Follows from Theorem 3.11 and the preceding remarks. ∎
4. Double cosets in $G$ and biequivariant sheaves on $(G,\omega)$
Fix a finite group scheme $G$, and a $3$-cocycle $\omega\in Z^{3}(G,\mathbb{G}_{m})$.
4.1. Double cosets in $G$
For the reader’s convenience, we first recall [GS, §4].
Let $\iota_{H}:H\xrightarrow{1:1}G$ and $\iota_{K}:K\xrightarrow{1:1}G$ be two embeddings of finite group schemes, and consider the right action of $H\times K$ on $G$ given by
(4.1) $\mu_{G\times H\times K}:G\times(H\times K)\to G,\,\,\,(g,h,k)\mapsto h^{-1}gk.$
The algebra map $\mu_{G\times H\times K}^{\sharp}:\mathscr{O}(G)\to\mathscr{O}(G)\otimes\mathscr{O}(H\times K)$ is given by
(4.2) $\mu_{G\times H\times K}^{\sharp}(f)=f_{2}\otimes\iota_{H}^{\sharp}{\rm S}(f_{1})\otimes\iota_{K}^{\sharp}(f_{3});\,\,\,f\in\mathscr{O}(G).$
Recall that since $H\times K$ is a finite group scheme, there exists a geometrical quotient finite scheme $Y:=G/(H\times K)$ (see §2.4).
For any closed point $g\in G(k)$, let $Z_{g}:=HgK\subset G$ denote the orbit of $g$ under the action (4.1). It is clear that $Z_{g}\in Y(k)$, and all closed points of $Y$ are such. As in §2.2), we have a direct sum decomposition
(4.3) ${\rm Coh}(Y)=\bigoplus_{Z_{g}\in Y(k)}{\rm Coh}(Y)_{Z_{g}}.$
For the remaining of this section, we fix $Z\in Y(k)$ with representative $g\in Z(k)$. Let $L^{g}:=H\cap gKg^{-1}=H\times_{G}\,gKg^{-1}$ denote the (group scheme-theoretical) intersection of $H$ and $gKg^{-1}$. Let $\iota_{g}:L^{g}\hookrightarrow G$ be the inclusion morphism, and consider the group scheme embedding of $L^{g}$ in $H\times K$ given by
(4.4) $\partial_{g}:=\left(\iota_{g},\left({\rm Ad}g^{-1}\right)\circ\iota_{g}\right):L^{g}\xrightarrow{1:1}H\times K,\;\;l\mapsto(l,g^{-1}lg).$
It is clear that the subgroup scheme $\partial_{g}(L^{g})\subset H\times K$ is the stabilizer of $g$ for the action (4.1). Namely, $Z$ is a quotient scheme for the free left action of $L^{g}$ on $H\times K$ given by
(4.5) $\mu_{L^{g}\times H\times K}:L^{g}\times(H\times K)\to H\times K,\,\,\,(l,h,k)\mapsto(lh,g^{-1}lgk).$
Thus, we have scheme isomorphisms
(4.6) $\begin{split}&\mathfrak{j}_{g}:L^{g}\backslash(H\times K)\xrightarrow{\cong}Z_{g},\,\,\,\overline{(h,k)}\mapsto h^{-1}gk,\,\,\,{\rm and}\\ &\mathfrak{j}_{g}^{-1}:Z_{g}\xrightarrow{\cong}L^{g}\backslash(H\times K),\,\,\,hgk\mapsto\overline{(h^{-1},k)}.\end{split}$
4.2. Biequivariant sheaves on $(G,\omega)$
Retain the setup of §4.1. For any $g\in G(k)$, define the $2$-cochain
(4.7) ${\rm w}_{g}:=\omega_{1}^{g^{-1}}\cdot\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right)\cdot\omega_{3}\in C^{2}(G,\mathbb{G}_{m})$
(see (2.13)). Also, let $p_{1}:H\times K\twoheadrightarrow H$ and $p_{2}:H\times K\twoheadrightarrow K$ be the obvious projection morphisms, and define the $2$-cochain
(4.8) $\begin{split}&{\rm W}_{g}:=\omega(g,g^{-1},g)\cdot p_{1}^{\sharp}(\omega_{3}^{1})p_{2}^{\sharp}(\omega_{1}^{1})p_{1}^{\sharp}(\omega_{2}^{-1})\otimes p_{1}^{\sharp}(\omega_{3}^{2})p_{2}^{\sharp}(\omega_{1}^{2})p_{2}^{\sharp}(\omega_{2}^{-2})\\ &=p_{1}^{\sharp\otimes 2}(\omega_{3})\cdot p_{2}^{\sharp\otimes 2}(\omega_{1})\cdot(p_{1}^{\sharp}\otimes p_{2}^{\sharp})(\omega_{2}^{-1})\in C^{2}(H\times K,\mathbb{G}_{m}).\end{split}$
Assume that $\psi\in C^{2}(H,\mathbb{G}_{m})$ and $\eta\in C^{2}(K,\mathbb{G}_{m})$ are such that $d\psi=\iota_{H}^{\sharp\otimes 3}(\omega)$ and $d\eta=\iota_{K}^{\sharp\otimes 3}(\omega)$. Set $\Psi:=\psi^{-1}\times\eta\in C^{2}(H\times K,\mathbb{G}_{m})$, and $\Theta_{g}:=\Psi^{-1}{\rm W}_{g}\in C^{2}(H\times K,\mathbb{G}_{m})$.
Recall the scheme embedding $\partial_{g}$ (4.4).
Lemma 4.1.
For any $Z\in Y(k)$ with representative $g\in Z(k)$, the following hold:
  1. (1)
    $\partial_{g}^{\sharp\otimes 2}\left(\Psi\right)=\psi^{-1}\cdot\eta^{g^{-1}}\in C^{2}(L^{g},\mathbb{G}_{m})$.
  2. (2)
    $\partial_{g}^{\sharp\otimes 2}\left({\rm W}_{g}\right)=\iota_{g}^{\sharp\otimes 2}\left({\rm w}_{g}\right)\in C^{2}(L^{g},\mathbb{G}_{m})$.
  3. (3)
    $\xi_{g}:=\partial_{g}^{\sharp\otimes 2}\left(\Theta_{g}\right)\in Z^{2}(L^{g},\mathbb{G}_{m})$ is a $2$-cocycle.
  4. (4)
    $(\partial_{g}^{\sharp\otimes 2}\otimes{\rm id})\left(d\Theta_{g}\right)=(\iota_{g}^{\sharp\otimes 2}\otimes p_{1}^{\sharp}\rho_{g})(\omega)\cdot(\iota_{g}^{\sharp\otimes 2}\otimes p_{2}^{\sharp}\lambda_{g^{-1}})(\omega^{-1})$.
Proof.
(1)-(2) Straightforward.
(3) We have to show that $d\xi_{g}=1$, i.e., $d\left(\psi^{-1}\cdot\eta^{g^{-1}}\right)=d{\rm w}_{g}$. Indeed, on the one hand, we have
$d\left(\psi^{-1}\cdot\eta^{g^{-1}}\right)=\left(d\psi^{-1}\right)\cdot d\left(\eta^{g^{-1}}\right)=\omega^{-1}\cdot\omega^{g^{-1}}.$
On the other hand, we have
$\begin{split}&d{\rm w}_{g}=\left(1\otimes{\rm w}_{g}\right)\cdot\left(\Delta\otimes{\rm id}\right)\left({\rm w}_{g}^{-1}\right)\cdot\left({\rm id}\otimes\Delta\right)\left({\rm w}_{g}\right)\cdot\left({\rm w}_{g}^{-1}\otimes 1\right)\\ &=\left(1\otimes\omega_{3}\right)\cdot\left(\Delta\otimes{\rm id}\right)\left(\omega_{3}^{-1}\right)\cdot\left({\rm id}\otimes\Delta\right)\left(\omega_{3}\right)\\ &\cdot\left(\Delta\otimes{\rm id}\right)\left(\omega_{1}^{g^{-1}}\right)^{-1}\cdot\left({\rm id}\otimes\Delta\right)\left(\omega_{1}^{g^{-1}}\right)\cdot\left(\omega_{1}^{g^{-1}}\otimes 1\right)^{-1}\\ &\cdot\left(\Delta\otimes{\rm id}\right)\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}\right)\cdot\left(1\otimes\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right)\right)\cdot\left(\omega_{3}^{-1}\otimes 1\right)\\ &\cdot\left(1\otimes\omega_{1}^{g^{-1}}\right)\cdot\left(\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}\right)\otimes 1\right)\cdot\left({\rm id}\otimes\Delta\right)\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right).\end{split}$
Now, using (2.12), it is straightforward to verify that
$\left(1\otimes\omega_{3}\right)\cdot\left(\Delta\otimes{\rm id}\right)\left(\omega_{3}^{-1}\right)\cdot\left({\rm id}\otimes\Delta\right)\left(\omega_{3}\right)=\left({\rm id}^{\otimes 2}\otimes\rho_{g}\right)\left(\omega\right)\cdot\omega^{-1};$
$\left(\Delta\otimes{\rm id}\right)\left(\omega_{1}^{g^{-1}}\right)^{-1}\cdot\left({\rm id}\otimes\Delta\right)\left(\omega_{1}^{g^{-1}}\right)=\omega^{g^{-1}}\cdot\left(\lambda_{g^{-1}}\otimes{\rm id}^{\otimes 2}\right)\left(\omega^{g^{-1}}\right)^{-1};$
$\left(\Delta\otimes{\rm id}\right)\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}\right)\cdot\left(1\otimes\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right)\right)\cdot\left(\omega_{3}^{-1}\otimes 1\right)$
$=\left({\rm id}\otimes\rho_{g}\otimes{\rm Ad}g^{-1}\right)\left(\omega\right)\cdot\left({\rm id}^{\otimes 2}\otimes\rho_{g}\right)\left(\omega^{-1}\right);$
$\left(1\otimes\omega_{1}^{g^{-1}}\right)\cdot\left(\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}\right)\otimes 1\right)\cdot\left({\rm id}\otimes\Delta\right)\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right)$
$=\left(\lambda_{g^{-1}}\otimes{\rm id}^{\otimes 2}\right)\left(\omega^{g^{-1}}\right)\cdot\left({\rm id}\otimes\rho_{g}\otimes{\rm Ad}g^{-1}\right)\left(\omega^{-1}\right).$
Thus, $d{\rm w}_{g}=\omega^{-1}\cdot\omega^{g^{-1}}$, as desired.
(4) Similar to (3). ∎
Now for any closed point $Z\in Y(k)$ with representative $g\in Z(k)$, choose a cleaving map (2.29) $\mathfrak{c}_{g}:\mathscr{O}(L^{g})\xrightarrow{1:1}\mathscr{O}(H\times K)$, and let
(4.9) $\alpha_{g}:\mathscr{O}(H\times K)\twoheadrightarrow\mathscr{O}(L^{g}\backslash(H\times K)),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}_{g}^{-1}\left(\partial_{g}^{\sharp}\left({\rm f}_{2}\right)\right)$
(see (2.32)). Let $\iota_{Z}:Z\hookrightarrow G$ be the inclusion morphism. Set
(4.10) ${\rm Ind}_{Z}:={\rm Ind}_{(L^{g},\xi_{g}^{-1})}^{\left((H,\psi),(K,\eta)\right)},\,\,\,{\rm F}_{Z}:={\rm F},\,\,\,\omega_{Z}:=\iota_{Z}^{\sharp\otimes 3}(\omega)$
(see Theorems 3.11, 3.13), and
(4.11) $\Phi_{g}:=\left(\iota_{Z}\circ\mathfrak{j}_{g}\right)^{\sharp\otimes 3}(\omega)=\mathfrak{j}_{g}^{\sharp\otimes 3}(\omega_{Z})\in C^{3}(L^{g}\backslash(H\times K),\mathbb{G}_{m}).$
Corollary 4.2.
For any closed point $Z\in Y(k)$ with representative $g\in Z(k)$, we have the following equivalences of abelian categories:
  1. (1)
    ${\rm Ind}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g}),$
    $(V,\ell)\mapsto\left(V\otimes^{\mathscr{O}(L^{g})_{\xi_{g}}}\left(\mathscr{O}(H\times K)\right),\lambda^{(\ell,\psi)}_{1},\rho_{\eta}\otimes_{k}{\rm id}_{V}\right)$.
  2. (2)
    ${\rm F}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g}),$
    $V\mapsto(\mathscr{O}(L^{g}\backslash(H\times K))\otimes_{k}V,\lambda_{V},\rho_{V})$.
  3. (3)
    $\mathfrak{j}_{g*}:{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g})\xrightarrow{\simeq}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(Z,\omega_{Z})$.
Proof.
Follows from Theorems 3.11, 3.13, and Lemma 4.1 for $A:=L^{g}$, $\Phi:=\Phi_{g}$, $(B,\beta):=(H,\psi)$, $(C,\gamma):=(K,\eta)$, $\partial:=\partial_{g}$, and $\xi:=\xi_{g}$. ∎
5. Module categories over ${\mathscr{C}}\left(G,\omega,H,\psi\right)$
Fix a finite group scheme-theoretical category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$ as in §2.12, and fix an indecomposable exact left $\mathscr{C}$}-module category
${\mathscr{M}}:=\mathscr{M}\left((H,\psi),(K,\eta)\right)={\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(G,\omega).$
In this section we apply Theorems 3.11, 3.13 and Corollary 4.2 to study the abelian structure of $\mathscr{M}$}.
5.1. The simple objects of $\mathscr{M}\left((H,\psi),(K,\eta)\right)$
Recall (§2.11) that $\psi\in C^{2}(H,\mathbb{G}_{m})$ and $\eta\in C^{2}(K,\mathbb{G}_{m})$ are such that $d\psi=\iota_{H}^{\sharp\otimes 3}(\omega)$ and $d\eta=\iota_{K}^{\sharp\otimes 3}(\omega)$.
Recall that $Y=G/(H\times K)$4.1). Fix $Z\in Y(k)$, and let
(5.1) ${\mathscr{M}}_{Z}:=\mathscr{M}_{Z}\left((H,\psi),(K,\eta)\right)\subset{\mathscr{M}}$
denote the full abelian subcategory of $\mathscr{M}$} consisting of all objects annihilated by the defining ideal $\mathscr{I}(Z)\subset\mathscr{O}(G)$ of $Z$. Namely, the inclusion morphism $\iota_{Z}:Z\hookrightarrow G$ is $(H,K)$-biequivariant, and ${\mathscr{M}}_{Z}$ is the image of the injective functor
(5.2) $\iota_{Z*}:{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(Z,\omega_{Z})\xrightarrow{1:1}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(G,\omega).$
(See Proposition 2.8.) Also, let $\overline{{\mathscr{M}}_{Z}}\subset{\mathscr{M}}$ denote the Serre closure of ${\mathscr{M}}_{Z}$ inside $\mathscr{M}$}, i.e., $\overline{{\mathscr{M}}_{Z}}$ is the full abelian subcategory of $\mathscr{M}$} consisting of all objects whose composition factors lie in ${\mathscr{M}}_{Z}$.
Fix a representative $g\in Z(k)$, and define the functor
(5.3) $\mathbf{Ind}_{Z}:=\iota_{Z*}\circ\mathfrak{j}_{g*}\circ{\rm Ind}_{Z}.$
Theorem 5.1.
Let ${\mathscr{M}}={\mathscr{M}}\left((H,\psi),(K,\eta)\right)$ be as above.
  1. (1)
    For any closed point $Z\in Y(k)$ with representative $g\in Z(k)$, we have an equivalence of abelian categories
    $\mathbf{Ind}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z},$
    $(V,\ell)\mapsto\left(\iota_{Z}\mathfrak{j}_{g}\right)_{*}\left(V\otimes^{\mathscr{O}(L^{g})_{\xi_{g}}}\mathscr{O}(H\times K),\lambda^{(\ell,\psi)}_{1},\rho_{\eta}\otimes_{k}{\rm id}_{V}\right).$
  2. (2)
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$ is simple, and simple objects of $\mathscr{M}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$.
  3. (3)
    We have a direct sum decomposition of abelian categories
    ${\mathscr{M}}=\bigoplus_{Z\in Y(k)}\overline{{\mathscr{M}}_{Z}}.$
Proof.
(1) Follows from Corollary 4.2.
(2) Assume $S\in{\mathscr{M}}$ is simple. Since for any $Z\in Y(k)$, the ideal $\mathscr{I}(Z)$ is $(H,K)$-bistable, it follows from Proposition 2.14(2) that either $\mathscr{I}(Z)S=0$ or $\mathscr{I}(Z)S=S$. If $\mathscr{I}(Z)S=S$ for every $Z\in Y(k)$, then $(\Pi_{Z}\mathscr{I}(Z))S=S$. But, $\Pi_{Z}\mathscr{I}(Z)$ is a nilpotent ideal of $\mathscr{O}(G)$ (being contained in the radical of $\mathscr{O}(G)$), so $S=0$, a contradiction.
Thus, there exists $Z\in Y(k)$ such that $\mathscr{I}(Z)S=0$. Assume that $\mathscr{I}(Z^{\prime})S=0$ for some $^{\prime}\in Y(k), ZZ$$^{\prime}\neq Z$. Then $\mathscr{O}(G)S=0$ (since $\mathscr{O}(G)=\mathscr{I}(Z)+\mathscr{I}(Z^{\prime})$), a contradiction. It follows that there exists a unique $Z\in Y(k)$ such that $\mathscr{I}(Z)S=0$, i.e., $S\in{\mathscr{M}}_{Z}$, so the claim follows from (1).
(3) Follows from (2), and the fact that there are no nontrivial cross extensions of $\mathscr{O}(G)$-modules which are supported on distinct closed points of $G$. ∎
5.2. Projectives in $\mathscr{M}\left((H,\psi),(K,\eta)\right)$
Let $\mathscr{M}$} be as in §5.1.
For any $Z\in Y(k)$ with representative $g\in Z(k)$, recall the functor ${\rm F}_{Z}$ from Corollary 4.2, and define the functor
(5.4) $\mathbf{F}_{Z}:=\iota_{Z*}\circ\mathfrak{j}_{g*}\circ{\rm F}_{Z}.$
Also, set $Z^{\circ}:=H^{\circ}gK^{\circ}$,
(5.5) $|Z^{\circ}|:=\frac{|H^{\circ}||K^{\circ}|}{|(L^{g})^{\circ}|}\in\mathbb{Z}^{\geq 1},\,\,\,{\rm and}\,\,\,|Z(k)|:=\frac{|H(k)||K(k)|}{|(L^{g})(k)|}\in\mathbb{Z}^{\geq 1}.$
Note that (2.3) induces a split exact sequence of schemes
$1\to Z^{\circ}\xrightarrow{i_{Z^{\circ}}}Z\mathrel{\mathop{\rightleftarrows}^{\pi_{Z}}_{q_{Z}}}Z(k)\to 1.$
Proposition 5.2.
For any $Z\in Y(k)$ with representative $g\in Z(k)$, the following hold:
  1. (1)
    The surjective $\mathscr{O}(G)$-linear algebra map
    $\chi_{Z}:=({\rm id}\otimes q_{Z}^{\sharp}):\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z)\twoheadrightarrow\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))$
    splits via the map
    $\nu_{Z}:=({\rm id}\otimes\pi_{Z}^{\sharp}):\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\xrightarrow{1:1}\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z).$
    Thus, $\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))$ is a direct summand of $\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z)$ as an $\mathscr{O}(G)$-module.
  2. (2)
    For any simple $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, we have an $\mathscr{M}$}-isomorphism
    $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right)\cong\mathscr{O}(Z^{\circ})\otimes_{k}P_{{\mathscr{M}}}\left(\mathbf{F}_{Z}(V)\right).$
    Here, $\mathscr{O}(G)$ acts diagonally on the left hand side, and $\mathscr{O}(H)_{\psi}$ and $\mathscr{O}(K)_{\eta}$ coact on its second factor, while the right hand side is a direct sum of $|Z^{\circ}|$ copies of $P_{{\mathscr{M}}}\left(\mathbf{F}_{Z}(V)\right)$.
Proof.
(1) Follows from the preceding remarks.
(2) By Example 3.12, $\mathbf{F}_{Z}\left(\mathscr{O}(L^{g})_{\xi_{g}^{-1}}\right)\cong\mathscr{O}(H\times K)$ in $\mathscr{M}$}, so
$\mathscr{O}(G)\otimes\mathbf{F}_{Z}\left(\mathscr{O}(L^{g})_{\xi_{g}^{-1}}\right)\cong\mathscr{O}(G)\otimes\mathscr{O}(H\times K)$
is projective in $\mathscr{M}$} (where $\mathscr{O}(G)$ acts diagonally, and $\mathscr{O}(H)_{\psi}$ and $\mathscr{O}(K)_{\eta}$ coact on the second factor via $\Delta_{\psi}$ and $\Delta_{\eta}$). Since $\mathscr{O}(G^{\circ})$ is a direct summand of $\mathscr{O}(G)$, and $\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right)$ is a direct summand of $\mathbf{F}_{Z}\left(\mathscr{O}(L^{g})_{\xi_{g}^{-1}}\right)$, it follows that the object $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right)$ of $\mathscr{M}$} is a direct summand of $\mathscr{O}(G)\otimes\mathscr{O}(H\times K)$, hence projective in $\mathscr{M}$}. Thus, $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right)$ is projective in $\mathscr{M}$}.
Now, on the one hand, we have
${\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(\mathscr{O}(L^{g})_{\xi_{g}^{-1}}\right),\mathbf{F}_{Z}(V)\right)$
$=$ ${\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathscr{O}(H\times K),\mathbf{F}_{Z}(V)\right)$
$=$ ${\rm Hom}_{{\rm Coh}(G,\omega)}\left(\mathscr{O}(G^{\circ}),\mathscr{O}(Z)\right)\otimes_{k}V.$
On the other hand, note that for any simple $W\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, the objects $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right)$ and $\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right)$ have the same composition factors as $\mathscr{O}(G)$-modules. Hence, if $V$ and $W$ are nonisomorphic simples in ${\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, then
${\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right),\mathbf{F}_{Z}(V)\right)=0.$
This implies that
${\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(\mathscr{O}(L^{g})_{\xi_{g}^{-1}}\right),\mathbf{F}_{Z}(V)\right)$
$=$ $\bigoplus_{W\in\mathcal{O}\left(\mathscr{O}(L^{g})_{\xi_{g}^{-1}}\right)}{\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right),\mathbf{F}_{Z}(V)\right)\otimes_{k}W$
$=$ ${\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right),\mathbf{F}_{Z}(V)\right)\otimes_{k}V.$
Thus, it follows from the above that
$\dim{\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right),\mathbf{F}_{Z}(V)\right)$
$=$ $\dim{\rm Hom}_{{\rm Coh}(G,\omega)}\left(\mathscr{O}\left(G^{\circ}\right),\mathscr{O}\left(Z\right)\right)=|Z^{\circ}|,$
which implies the statement. ∎
Theorem 5.3.
Let ${\mathscr{M}}={\mathscr{M}}\left((H,\psi),(K,\eta)\right)$ be as above.
  1. (1)
    For any $Z\in Y(k)$ with representative $g\in Z(k)$, we have an equivalence of abelian categories
    $\mathbf{F}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z},\,\,\,V\mapsto\iota_{Z*}\left(\mathscr{O}(Z)\otimes_{k}V,\lambda^{g}_{V},\rho^{g}_{V}\right),$
    where
    $\lambda^{g}_{V}:=\left({\rm id}\otimes\left(\mathfrak{j}_{g}^{\sharp}\right)^{-1}\otimes{\rm id}\right)\lambda_{V}\left(\mathfrak{j}_{g}^{\sharp}\otimes{\rm id}\right),\,\,\rho^{g}_{V}:=\left(\left(\mathfrak{j}_{g}^{\sharp}\right)^{-1}\otimes{\rm id}^{\otimes 2}\right)\rho_{V}\left(\mathfrak{j}_{g}^{\sharp}\otimes{\rm id}\right)$
    (see (3.27)-(3.28)).
  2. (2)
    For any simple $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, we have
    $P_{{\mathscr{M}}}\left(\mathbf{F}_{Z}(V)\right)\cong\left(\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V),L^{g}_{V},R^{g}_{V}\right),$
    where $\mathscr{O}(G)$ acts diagonally and
    $L^{g}_{V}:=l^{g}\cdot\left({\rm id}\otimes\chi_{Z}\otimes{\rm id}\right)(12)\left({\rm id}\otimes\lambda^{g}_{P_{(L^{g},\xi_{g}^{-1})}(V)}\right)\left(\nu_{Z}\otimes{\rm id}\right),$
    $R^{g}_{V}:=r^{g}\cdot\left(\chi_{Z}\otimes{\rm id}\otimes{\rm id}\right)\left({\rm id}\otimes\rho^{g}_{P_{(L^{g},\xi_{g}^{-1})}(V)}\right)\left(\nu_{Z}\otimes{\rm id}\right),$
    $l^{g}:=(\iota_{H}^{\sharp}\otimes\iota_{G^{\circ}}^{\sharp}\otimes\iota_{Z}^{\sharp})(\omega^{-1})\cdot(\iota_{G^{\circ}}^{\sharp}\otimes\iota_{H}^{\sharp}\otimes\iota_{Z}^{\sharp})(\omega)$, and $r^{g}:=(\iota_{G^{\circ}}^{\sharp}\otimes\iota_{Z}^{\sharp}\otimes\iota_{K}^{\sharp})(\omega)$.
Proof.
(1) Follows from Theorem 3.13 and 5.1.
(2) We have $\mathscr{M}$}-isomorphisms
$\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(V)\right)=\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z)\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V)$
$\cong$ $\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}\mathscr{O}(Z^{\circ})\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V),$
thus by Proposition 5.2, $\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V)$ is a direct summand in a projective object of $\mathscr{M}$}, hence is projective. Moreover, by Proposition 5.2 again, we have
$\dim{\rm Hom}_{{\mathscr{M}}}\left(\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V),\mathbf{F}_{Z}(V)\right)=1.$
Hence, $\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}P_{(L^{g},\xi_{g}^{-1})}(V)$ is the projective cover of $\mathbf{F}_{Z}(V)$ in $\mathscr{M}$}, as claimed. ∎
Example 5.4.
[…]
$\mathbf{F}_{Hg}:{\rm Vec}\xrightarrow{\simeq}{\rm Coh}^{\left(\left(H,\psi\right),\left(1,1\right)\right)}(Hg),\,\,\,V\mapsto\mathscr{O}(Hg)\otimes_{k}V,$
is an equivalence of abelian categories (where $\mathscr{O}(H)$ acts on $\mathscr{O}(Hg)$ via multiplication by $\rho_{g^{-1}}(\cdot)$ and $\mathscr{O}(H)_{\psi}$-coacts via $\Delta_{\psi}$).
Now by Theorem 5.3, there is a bijection between the set of closed points $\bar{g}:=Hg\in Y(k)$ (i.e., isomorphism classes of simples of ${\rm Coh}(H\backslash G)$) and simple objects of $\mathscr{M}$, assigning $\bar{g}$ to $\mathscr{O}(Hg)\otimes_{k}k=\mathscr{O}(Hg)$, and we have $\overline{{\mathscr{M}}_{Hg}}=\langle\mathscr{O}(Hg)\rangle$.
Moreover, by Theorem 5.3 again, for any simple $\mathscr{O}(Hg)\in{\mathscr{M}}$,
$P_{{\mathscr{M}}}(\mathscr{O}(Hg))\cong\mathscr{O}(G^{\circ})\otimes\mathscr{O}(H(k)g)\otimes_{k}k\cong P_{Hg}\otimes_{k}k\cong P_{Hg}$
as $\mathscr{O}(G)$-modules, where $\mathscr{O}(H)_{\psi}$ coacts on $P_{Hg}\otimes_{k}k\cong P_{Hg}$ via the map $\lambda_{k}=\Delta_{\psi}$. (Compare with Remark 2.24.) ∎
5.3. Fiber functors on ${\mathscr{C}}(G,\omega,H,\psi)$
Recall that a fiber functor on a finite tensor category is the same as a module category of rank $1$.
The next corollary generalizes [GS, Corollary 5.8] (see [GS, Example 5.9] for concrete examples).
Corollary 5.5.
Let ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$ be a group scheme-theoretical category. There is a bijection between equivalence classes of fiber functors on $\mathscr{C}$} and equivalence classes of pairs $(K,\eta)$, where $K\subset G$ is a closed subgroup scheme and $\eta\in C^{2}(K,\mathbb{G}_{m})$, such that $d\eta=\iota_{K}^{\sharp\otimes 3}(\omega)$, $HK=G$, and $\xi^{-1}:=\xi_{1}^{-1}\in Z^{2}(H\cap K,\mathbb{G}_{m})$ is nondegenerate.
Proof.
Let ${\mathscr{M}}:=\mathscr{M}\left(\left(H,\psi\right),\left(K,\eta\right)\right)$ be an indecomposable exact module category over $\mathscr{C}$}. By Theorem 5.1, ${\mathscr{M}}\simeq{\rm Vec}$ if and only if ${\mathscr{M}}={\mathscr{M}}_{HK}$ and ${\rm Corep}_{k}(\mathscr{O}(H\cap K)_{\xi^{-1}})\simeq{\rm Vec}$. Thus, the statement follows from the fact that ${\mathscr{M}}={\mathscr{M}}_{HK}$ if and only if $\iota_{HK*}$ is an equivalence, i.e., if and only if $G=HK$. ∎
6. The structure of ${\mathscr{C}}\left(G,\omega,H,\psi\right)$
Fix a group scheme-theoretical category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$ §2.12. For any closed point $g\in G(k)$, let $H^{g}:=H\cap gHg^{-1}$. Note that $\xi_{1}=1$.
Recall that $Y=G/(H\times K)$4.1).
Theorem 6.1.
The following hold:
  1. (1)
    For any closed point $Z\in Y(k)$ with representative $g\in Z(k)$, we have an equivalence of abelian categories
    $\mathbf{Ind}_{Z}:{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}\mathscr{C}_{Z},$
    $(V,\ell)\mapsto(\iota_{Z}\mathfrak{j}_{g})_{*}\left(V\otimes^{\mathscr{O}(H^{g})_{\xi_{g}}}\mathscr{O}(H\times H),\lambda^{(\ell,\psi)}_{1},\rho_{\psi}\otimes_{k}{\rm id}_{V}\right)$.
    In particular,
    $\mathbf{Ind}_{H}:{\rm Rep}(H)_{k}\xrightarrow{\simeq}\mathscr{C}_{H},$
    $(V,\ell)\mapsto(\iota_{H}\mathfrak{j}_{1})_{*}\left(V\otimes^{\mathscr{O}(H)}\mathscr{O}(H\times H),\lambda^{(\ell,\psi)}_{1},\rho_{\psi}\otimes_{k}{\rm id}_{V}\right),$
    is an equivalence of tensor categories.
  2. (2)
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$ is simple, and simple objects of $\mathscr{C}$}, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$. Moreover, we have a direct sum decomposition of abelian categories
    ${\mathscr{C}}=\bigoplus_{Z\in Y(k)}\overline{{\mathscr{C}}_{Z}},$
    and $\overline{{\mathscr{C}}_{H}}\subset{\mathscr{C}}$ is a tensor subcategory.
  3. (3)
    For any $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have
    ${\rm FPdim}\left(\mathbf{Ind}_{Z}(V)\right)=\frac{|H|}{|H^{g}|}{\rm dim}(V).$
  4. (4)
    For any $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have $\mathbf{Ind}_{Z}(V)^{*}\cong\mathbf{Ind}_{Z^{-1}}(V^{*})$, where $Z^{-1}\in Y(k)$ such that $g^{-1}\in Z^{-1}(k)$.
Proof.
Follow from Theorem 5.1. ∎
Theorem 6.2.
The following hold:
  1. (1)
    For any closed point $Z\in Y(k)$ with representative $g\in Z(k)$, we have an equivalence of abelian categories
    $\mathbf{F}_{Z}:{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{C}}_{Z},\,\,\,V\mapsto\iota_{Z*}\left(\mathscr{O}(Z)\otimes_{k}V,\lambda^{g}_{V},\rho^{g}_{V}\right).$
    In particular, we have an equivalence of tensor categories
    $\mathbf{F}_{H}:{\rm Rep}(H)_{k}\xrightarrow{\simeq}\mathscr{C}_{H},\,\,\,V\mapsto\iota_{H*}\left(\mathscr{O}(H)\otimes_{k}V,\lambda^{1}_{V},\rho^{1}_{V}\right).$
  2. (2)
    For any $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have $\mathbf{F}_{Z}(V)^{*}\cong\mathbf{F}_{Z^{-1}}(V^{*})$.
  3. (3)
    For any simple $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have
    $P_{{\mathscr{C}}}\left(\mathbf{F}_{Z}(V)\right)\cong\left(\mathscr{O}(G^{\circ})\otimes\mathscr{O}(Z(k))\otimes_{k}P_{(H^{g},\xi_{g}^{-1})}(V),L^{g}_{V},R^{g}_{V}\right),$
    and $l^{g}:=(\iota_{H}^{\sharp}\otimes\iota_{G^{\circ}}^{\sharp}\otimes\iota_{Z}^{\sharp})(\omega^{-1})\cdot(\iota_{G^{\circ}}^{\sharp}\otimes\iota_{H}^{\sharp}\otimes\iota_{Z}^{\sharp})(\omega)$, and $r^{g}:=(\iota_{G^{\circ}}^{\sharp}\otimes\iota_{Z}^{\sharp}\otimes\iota_{H}^{\sharp})(\omega)$. In particular,
    ${\rm FPdim}\left(P_{{\mathscr{C}}}\left(\mathbf{F}_{Z}\left(V\right)\right)\right)=\frac{|G^{\circ}||H(k)|}{|H^{\circ}||H^{g}(k)|}{\rm dim}\left(P_{(H^{g},\xi_{g}^{-1})}\left(V\right)\right).$
  4. (4)
    For any closed point $Z\in Y(k)$, ${\rm FPdim}\left(\overline{{\mathscr{C}}_{Z}}\right)=|G^{\circ}||Z(k)|$.
  5. (5)
    If ${\rm Rep}_{k}(H)$ is unimodular, so is $\mathscr{C}$} (but not necessarily vice versa).
Proof.
(1)-(4) Follow from Theorem 5.3 in a straightforward manner.
(5) Recall that ${\rm Rep}_{k}(H)$ is unimodular if and only if $P(\mathbf{1})\cong P(\mathbf{1})^{*}$, where $P(\mathbf{1})$ is the projective cover of $\mathbf{1}$ in ${\rm Rep}_{k}(H)$. Thus, if ${\rm Rep}_{k}(H)$ is unimodular then
$|H^{\circ}|P_{{\mathscr{C}}}(\mathbf{1})\cong\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})\right)$
$\cong$ $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})^{*}\right)\cong\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})\right)^{*}$
$\cong$ $\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})\right)\right)^{*}\cong|H^{\circ}|P_{{\mathscr{C}}}(\mathbf{1})^{*},$
so $P_{{\mathscr{C}}}(\mathbf{1})\cong P_{{\mathscr{C}}}(\mathbf{1})^{*}$. ∎
Remark 6.3.
Theorem 6.2(5) for étale $G$ with $\omega=1$ follows from [Y].
6.1. The étale case
Assume $G$ is étale, i.e, $G=G(k)$. The following result is known in characteristic 0 [GN, O].
Corollary 6.4.
The following hold:
  1. (1)
    For any $(H,H)$-double coset $Z$ with representative $g\in G$, and simple $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have
    $P_{{\mathscr{C}}}\left(\mathbf{F}_{Z}(V)\right)=\mathbf{F}_{Z}\left(P_{(H^{g},\xi_{g}^{-1})}\left(V\right)\right)=\left(\mathscr{O}\left(Z\right)\otimes_{k}P_{(H^{g},\xi_{g}^{-1})}\left(V\right),L_{V}^{g},R^{g}_{V}\right),$
    and $l^{g}=r^{g}=1$.
  2. (2)
    We have a direct sum decomposition of abelian categories
    ${\mathscr{C}}\simeq\bigoplus_{Z\in Y}{\rm Rep}(H^{g},\xi_{g}^{-1})_{k},$
    i.e., $\overline{{\mathscr{C}}_{Z}}={\mathscr{C}}_{Z}$ for every $Z$.
  3. (3)
    $\mathscr{C}$} is fusion if and only if $p$ does not divide $|H|$.
Proof.
Follow immediately from Theorem 6.2. ∎
6.2. The connected case
Assume that $G=G^{\circ}$. Recall the group scheme embedding $\partial:=\partial_{1}:H\xrightarrow{1:1}H\times H$, $h\mapsto(h,h)$ (4.4). It is clear that $\partial^{\sharp}:\mathscr{O}(H)^{\otimes 2}\to\mathscr{O}(H)$ is the multiplication map of $\mathscr{O}(H)$.
Note that the map
$\mathfrak{c}:\mathscr{O}(H)\to\mathscr{O}(H\times H),\,\,\,f\mapsto 1\otimes f,$
is a cleaving map (2.29) with convolution inverse
$\mathfrak{c}^{-1}:\mathscr{O}(H)\to\mathscr{O}(H\times H),\,\,\,f\mapsto 1\otimes{\rm S}(f).$
Since the induced $2$-cocycle $\sigma$ (2.19) is trivial, it follows that
$\phi:\mathscr{O}\left(H\backslash\left(H\times H\right)\right)\otimes\mathscr{O}(H)\xrightarrow{\cong}\mathscr{O}(H\times H),\,\,\,{\rm f}\otimes f\mapsto f^{\prime}(1\otimes f),$
$\phi^{-1}:\mathscr{O}(H\times H)\xrightarrow{\cong}\mathscr{O}\left(H\backslash\left(H\times H\right)\right)\otimes\mathscr{O}(H),\,\,\,{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\partial^{\sharp}({\rm f}_{2}))\otimes\partial^{\sharp}({\rm f}_{3}),$
and $\alpha:\mathscr{O}(H\times H)\twoheadrightarrow\mathscr{O}\left(H\backslash\left(H\times H\right)\right)$, ${\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\partial^{\sharp}({\rm f}_{2}))$ (see §2.9).
For any $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, set ${\rm F}_{V}=\left({\rm id}\otimes\alpha\right)_{21}$ (see §3.4). We have
${\rm F}_{V}:V\otimes^{\mathscr{O}(H)}\mathscr{O}(H\times H)\xrightarrow{\cong}\mathscr{O}(H\backslash(H\times H))\otimes_{k}V,$
$v\otimes{\rm f}\mapsto{\rm f}_{1}\mathfrak{c}^{-1}(\partial^{\sharp}({\rm f}_{2}))\otimes v,$
and
${\rm F}_{V}^{-1}:\mathscr{O}(H\backslash(H\times H))\otimes_{k}V\xrightarrow{\cong}V\otimes^{\mathscr{O}(H)}\mathscr{O}(H\times H),$
${\rm f}\otimes v\mapsto v^{0}\otimes{\rm f}\mathfrak{c}(v^{-1}).$
Recall also the maps $\lambda_{\psi}$ and $\rho_{\psi}$ (3.17), and the maps
$\lambda_{V}:=({\rm id}\otimes{\rm F}_{V})(12)({\rm id}\otimes\lambda_{\psi}){\rm F}_{V}^{-1},\,\,\,{\rm and}\,\,\,\rho_{V}:=({\rm F}_{V}\otimes{\rm id})({\rm id}\otimes\rho_{\psi}){\rm F}_{V}^{-1}$
(see (3.27)-(3.28)). Next we give an explicit formula of these coactions.
Lemma 6.5.
For any $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, the following hold:
  1. (1)
    $\lambda_{V}:\mathscr{O}(H\backslash(H\times H))\otimes_{k}V\to\mathscr{O}(H)_{\psi}\otimes\mathscr{O}(H\backslash(H\times H))\otimes_{k}V$, ${\rm f}\otimes v\mapsto\psi^{1}j_{1}^{\sharp}({\rm f}_{1})\otimes{\rm f}_{2}\left(\psi^{2}_{1}\otimes\partial^{\sharp}{\rm S}({\rm f}_{3}){\rm S}(\psi^{2}_{2})\right)\otimes v$.
  2. (2)
    $\rho_{V}:\mathscr{O}(H\backslash(H\times H))\otimes_{k}V\to\mathscr{O}(H\backslash(H\times H))\otimes_{k}V\otimes\mathscr{O}(H)_{\psi}$, ${\rm f}\otimes v\mapsto{\rm f}_{1}(1\otimes\partial^{\sharp}{\rm S}({\rm f}_{2}))\otimes v^{0}\otimes j_{2}^{\sharp}({\rm f}_{3})v^{-1}$.
Proof.
(1) For every ${\rm f}\otimes v\in\mathscr{O}(H\backslash(H\times H))\otimes_{k}V$, we have
$\begin{split}&\lambda_{V}({\rm f}\otimes v)=({\rm id}\otimes{\rm F}_{V})(12)({\rm id}\otimes\lambda_{\psi}){\rm F}_{V}^{-1}({\rm f}\otimes v)\\ &=({\rm id}\otimes{\rm F}_{V})(12)({\rm id}\otimes\lambda_{\psi})(v^{0}\otimes{\rm f}\mathfrak{c}(v^{-1}))\\ &=({\rm id}\otimes{\rm F}_{V})(12)\left(v^{0}\otimes\psi^{1}j_{1}^{\sharp}({\rm f}_{1}\mathfrak{c}(v^{-1})_{1})\otimes(\psi^{2}\otimes 1){\rm f}_{2}\mathfrak{c}(v^{-1})_{2}\right)\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1}\mathfrak{c}(v^{-1})_{1})\otimes{\rm F}_{V}\left(v^{0}\otimes(\psi^{2}\otimes 1){\rm f}_{2}\mathfrak{c}(v^{-1})_{2}\right)\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1}\mathfrak{c}(v^{-1})_{1})\otimes(\psi^{2}_{1}\otimes 1){\rm f}_{2}\mathfrak{c}(v^{-1})_{2}\mathfrak{c}^{-1}\left(\partial^{\sharp}((\psi^{2}_{2}\otimes 1){\rm f}_{3}\mathfrak{c}(v^{-1})_{3})\right)\otimes v^{0}\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1}(1\otimes v^{-1}))\otimes{\rm f}_{2}(\psi^{2}_{1}\otimes v^{-2})\mathfrak{c}^{-1}\left(\partial^{\sharp}({\rm f}_{3}(\psi^{2}_{2}\otimes v^{-3}))\right)\otimes v^{0}\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1})\varepsilon(v^{-1})\otimes{\rm f}_{2}(\psi^{2}_{1}\otimes v^{-2})\left(1\otimes\partial^{\sharp}{\rm S}({\rm f}_{3}(\psi^{2}_{2}\otimes v^{-3}))\right)\otimes v^{0}\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1})\otimes{\rm f}_{2}\left(\psi^{2}_{1}\otimes v^{-1}\partial^{\sharp}{\rm S}({\rm f}_{3}(\psi^{2}_{2}\otimes v^{-2}))\right)\otimes v^{0}\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1})\otimes{\rm f}_{2}\left(\psi^{2}_{1}\otimes v^{-1}\partial^{\sharp}{\rm S}({\rm f}_{3}){\rm S}(\psi^{2}_{2}){\rm S}(v^{-2})\right)\otimes v^{0}\\ &=\psi^{1}j_{1}^{\sharp}({\rm f}_{1})\otimes{\rm f}_{2}\left(\psi^{2}_{1}\otimes\partial^{\sharp}{\rm S}({\rm f}_{3}){\rm S}(\psi^{2}_{2})\right)\otimes v,\end{split}$
as claimed.
(2) For every ${\rm f}\otimes v\in\mathscr{O}(H\backslash(H\times H))\otimes_{k}V$, we have
$\begin{split}&\rho_{V}({\rm f}\otimes v)=({\rm F}_{V}\otimes{\rm id})({\rm id}\otimes\rho_{\psi}){\rm F}_{V}^{-1}({\rm f}\otimes v)\\ &=({\rm F}_{V}\otimes{\rm id})({\rm id}\otimes\rho_{\psi})(v^{0}\otimes{\rm f}\mathfrak{c}(v^{-1}))\\ &=({\rm F}_{V}\otimes{\rm id})\left(v^{0}\otimes(1\otimes\psi^{1}){\rm f}_{1}\mathfrak{c}(v^{-1})_{1}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{2}\mathfrak{c}(v^{-1})_{2})\right)\\ &={\rm F}_{V}\left(v^{0}\otimes(1\otimes\psi^{1}){\rm f}_{1}\mathfrak{c}(v^{-1})_{1}\right)\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{2}\mathfrak{c}(v^{-1})_{2})\\ &=(1\otimes\psi^{1}_{1}){\rm f}_{1}\mathfrak{c}(v^{-1})_{1}\mathfrak{c}^{-1}(\partial^{\sharp}((1\otimes\psi^{1}_{2}){\rm f}_{2}\mathfrak{c}(v^{-1})_{2}))\otimes v^{0}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{3}\mathfrak{c}(v^{-1})_{3})\\ &={\rm f}_{1}(1\otimes\psi^{1}_{1}v^{-1})(1\otimes{\rm S}(\psi^{1}_{2})\partial^{\sharp}{\rm S}({\rm f}_{2}\mathfrak{c}(v^{-1})_{2}))\otimes v^{0}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{3}\mathfrak{c}(v^{-1})_{3})\\ &={\rm f}_{1}(1\otimes\psi^{1}_{1}v^{-1})(1\otimes{\rm S}(\psi^{1}_{2})\partial^{\sharp}{\rm S}({\rm f}_{2}(1\otimes v^{-2})))\otimes v^{0}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{3}(1\otimes v^{-3}))\\ &={\rm f}_{1}(1\otimes\psi^{1}_{1}v^{-1})(1\otimes{\rm S}(\psi^{1}_{2})\partial^{\sharp}{\rm S}({\rm f}_{2}(1\otimes v^{-2})))\otimes v^{0}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{3})v^{-3}\\ &={\rm f}_{1}(1\otimes\psi^{1}_{1}v^{-1}{\rm S}(\psi^{1}_{2})\partial^{\sharp}{\rm S}({\rm f}_{2}){\rm S}(v^{-2}))\otimes v^{0}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{3})v^{-3}\\ &={\rm f}_{1}(1\otimes\psi^{1}_{1}{\rm S}(\psi^{1}_{2})\partial^{\sharp}{\rm S}({\rm f}_{2}))\otimes v^{0}\otimes\psi^{2}j_{2}^{\sharp}({\rm f}_{3})v^{-1}\\ &={\rm f}_{1}(1\otimes\partial^{\sharp}{\rm S}({\rm f}_{2}))\otimes v^{0}\otimes j_{2}^{\sharp}({\rm f}_{3})v^{-1},\end{split}$
as claimed. ∎
Now recall the scheme isomorphism $\mathfrak{j}:=\mathfrak{j}_{1}$ (4.6). We have
$\mathfrak{j}^{\sharp}:\mathscr{O}(H)\xrightarrow{\cong}\mathscr{O}\left(H\backslash\left(H\times H\right)\right),\,\,\,f\mapsto{\rm S}(f_{1})\otimes f_{2},$
and
$(\mathfrak{j}^{\sharp})^{-1}:\mathscr{O}\left(H\backslash\left(H\times H\right)\right)\xrightarrow{\cong}\mathscr{O}(H),\,\,\,{\rm f}\mapsto(\varepsilon\bar{\otimes}{\rm id})({\rm f}).$
Theorem 6.6.
The following hold:
  1. (1)
    We have ${\mathscr{C}}=\overline{{\mathscr{C}}_{H}}$.
  2. (2)
    We have an equivalence of tensor categories
    $\mathbf{F}:=\mathbf{F}_{H}:{\rm Corep}_{k}(\mathscr{O}(H))\xrightarrow{\simeq}{\mathscr{C}}_{H},\,\,\,V\mapsto\iota_{*}\left(\mathscr{O}(H)\otimes_{k}V,\lambda^{1}_{V},\rho^{1}_{V}\right),$
    where for every $f\otimes v\in\mathscr{O}(H)\otimes V$,
    $\lambda^{1}_{V}\left(f\otimes v\right)=\psi^{1}{\rm S}(f_{2})\otimes{\rm S}(\psi^{2})f_{1}\otimes v,\,\,\,{\rm and}\,\,\,\rho^{1}_{V}\left(f\otimes v\right)=f_{1}\otimes v^{0}\otimes f_{2}v^{-1}.$
  3. (3)
    For any simple $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, we have
    $P_{{\mathscr{C}}}\left(\mathbf{F}(V)\right)\cong\left(\mathscr{O}(G)\otimes_{k}P_{H}(V),L^{1}_{V},R^{1}_{V}\right),$
    where $L^{1}_{V}$ is trivial, $R^{1}_{V}={\rm id}\otimes\rho^{1}_{P_{G}(V)}$, and $l^{1}=r^{1}=1$.
  4. (4)
    For any $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, $\text{\rm FPdim}\left(P_{{\mathscr{C}}}\left(\mathbf{F}\left(V\right)\right)\right)=\frac{|G|}{|H|}\dim\left(P_{H}\left(V\right)\right)$.
Proof.
Follow from Theorem 6.2, except for the formulas of $\lambda^{1}_{V}$, $\rho^{1}_{V}$, $L^{1}_{V}$ and $R^{1}_{V}$.
(2) By Lemma 6.5, for every $f\otimes v\in\mathscr{O}(H)\otimes V$, we have
$\lambda_{V}\left(\mathfrak{j}^{\sharp}\otimes{\rm id}\right)\left(f\otimes v\right)=\lambda_{V}\left({\rm S}(f_{1})\otimes f_{2}\otimes v\right)$
$=$ $\psi^{1}j_{1}^{\sharp}({\rm S}(f_{1})\otimes f_{2})_{1}\otimes({\rm S}(f_{1})\otimes f_{2})_{2}\left(\psi^{2}_{1}\otimes\partial^{\sharp}{\rm S}({\rm S}(f_{1})\otimes f_{2})_{3}{\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}j_{1}^{\sharp}({\rm S}(f_{1})_{1}\otimes f_{2})\otimes({\rm S}(f_{1})_{2}\otimes f_{3})\left(\psi^{2}_{1}\otimes\partial^{\sharp}{\rm S}({\rm S}(f_{1})_{3}\otimes f_{4}){\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}j_{1}^{\sharp}({\rm S}(f_{3})\otimes f_{4})\otimes({\rm S}(f_{2})\otimes f_{5})\left(\psi^{2}_{1}\otimes\partial^{\sharp}{\rm S}({\rm S}(f_{1})\otimes f_{6}){\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}{\rm S}(f_{3})\varepsilon(f_{4})\otimes({\rm S}(f_{2})\otimes f_{5})\left(\psi^{2}_{1}\otimes\partial^{\sharp}(f_{1}\otimes{\rm S}(f_{6})){\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}{\rm S}(f_{3})\varepsilon(f_{4})\otimes({\rm S}(f_{2})\otimes f_{5})\left(\psi^{2}_{1}\otimes f_{1}{\rm S}(f_{6}){\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}{\rm S}(f_{3})\otimes({\rm S}(f_{2})\otimes f_{4})\left(\psi^{2}_{1}\otimes f_{1}{\rm S}(f_{5}){\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}{\rm S}(f_{3})\otimes\left({\rm S}(f_{2})\psi^{2}_{1}\otimes f_{1}f_{4}{\rm S}(f_{5}){\rm S}(\psi^{2}_{2})\right)\otimes v$
$=$ $\psi^{1}{\rm S}(f_{3})\otimes{\rm S}(f_{2})\psi^{2}_{1}\otimes f_{1}\varepsilon(f_{4}){\rm S}(\psi^{2}_{2})\otimes v$
$=$ $\psi^{1}{\rm S}(f_{3})\otimes{\rm S}(f_{2})\psi^{2}_{1}\otimes f_{1}{\rm S}(\psi^{2}_{2})\otimes v.$
Thus, it follows that
$\lambda^{1}_{V}\left(f\otimes v\right)=\left({\rm id}\otimes\left(\mathfrak{j}^{\sharp}\right)^{-1}\otimes{\rm id}\right)\lambda_{V}\left(\mathfrak{j}^{\sharp}\otimes{\rm id}\right)\left(f\otimes v\right)$
$=$ $\left({\rm id}\otimes\left(\mathfrak{j}^{\sharp}\right)^{-1}\otimes{\rm id}\right)\left(\psi^{1}{\rm S}(f_{3})\otimes{\rm S}(f_{2})\psi^{2}_{1}\otimes f_{1}{\rm S}(\psi^{2}_{2})\otimes v\right)$
$=$ $\psi^{1}{\rm S}(f_{3})\otimes\varepsilon\left({\rm S}(f_{2})\psi^{2}_{1}\right)f_{1}{\rm S}(\psi^{2}_{2})\otimes v=\psi^{1}{\rm S}(f_{2})\otimes{\rm S}(\psi^{2})f_{1}\otimes v,$
as claimed.
To compute $\rho_{V}$, by Lemma 6.5, for every $f\otimes v\in\mathscr{O}(H)\otimes V$, we have
$\rho_{V}\left(\mathfrak{j}^{\sharp}\otimes{\rm id}\right)\left(f\otimes v\right)=\rho_{V}\left({\rm S}\left(f_{1}\right)\otimes f_{2}\otimes v\right)$
$=$ $\left({\rm S}\left(f_{1}\right)\otimes f_{2}\right)_{1}\left(1\otimes\partial^{\sharp}{\rm S}\left({\rm S}\left(f_{1}\right)\otimes f_{2}\right)_{2}\right)\otimes v^{0}\otimes j_{2}^{\sharp}\left({\rm S}\left(f_{1}\right)\otimes f_{2}\right)_{3}v^{-1}$
$=$ $({\rm S}(f_{3})\otimes f_{4})(1\otimes\partial^{\sharp}{\rm S}({\rm S}(f_{2})\otimes f_{5}))\otimes v^{0}\otimes j_{2}^{\sharp}({\rm S}(f_{1})\otimes f_{6})v^{-1}$
$=$ $({\rm S}(f_{3})\otimes f_{4})(1\otimes\partial^{\sharp}\left(f_{2}\otimes{\rm S}(f_{5})\right)\otimes v^{0}\otimes\varepsilon{\rm S}(f_{1})f_{6}v^{-1}$
$=$ $({\rm S}(f_{2})\otimes f_{3})\left(1\otimes f_{1}{\rm S}(f_{4})\right)\otimes v^{0}\otimes f_{5}v^{-1}$
$=$ ${\rm S}(f_{2})\otimes f_{1}f_{3}{\rm S}(f_{4})\otimes v^{0}\otimes f_{5}v^{-1}={\rm S}(f_{2})\otimes f_{1}\otimes v^{0}\otimes f_{3}v^{-1}.$
Thus, it follows that
$\rho^{1}_{V}\left(f\otimes v\right)=\left(\left(\mathfrak{j}^{\sharp}\right)^{-1}\otimes{\rm id}^{\otimes 2}\right)\rho_{V}\left(\mathfrak{j}^{\sharp}\otimes{\rm id}\right)\left(f\otimes v\right)$
$=$ $\left(\left(\mathfrak{j}^{\sharp}\right)^{-1}\otimes{\rm id}^{\otimes 2}\right)\left({\rm S}(f_{2})\otimes f_{1}\otimes v^{0}\otimes f_{3}v^{-1}\right)$
$=$ $\varepsilon{\rm S}(f_{2})f_{1}\otimes v^{0}\otimes f_{3}v^{-1}=f_{1}\otimes v^{0}\otimes f_{2}v^{-1},$
as claimed.
(3) By (1), (2), and Proposition 5.2, we have
$\lambda^{1}_{P_{H}(V)}\left(1\otimes x\right)=\psi^{1}\otimes{\rm S}(\psi^{2})\otimes x;\,\,\,x\in P_{H}(V),$
$\rho^{1}_{P_{H}(V)}\left(1\otimes x\right)=1\otimes x^{0}\otimes x^{-1};\,\,\,x\in P_{H}(V),$
$\chi:=\chi_{H}:\mathscr{O}(G)\otimes\mathscr{O}(H)\twoheadrightarrow\mathscr{O}(G),\,\,\,{\bf f}\otimes f\mapsto{\bf f}\varepsilon(f),$
${\rm and}\,\,\,\nu:=\nu_{H}:\mathscr{O}(G)\xrightarrow{1:1}\mathscr{O}(G)\otimes\mathscr{O}(H),\,\,\,{\bf f}\mapsto{\bf f}\otimes 1.$
Thus, by Theorem 5.3, for any ${\bf f}\otimes x\in\mathscr{O}(G)\otimes_{k}P_{H}(V)$, we have
$\begin{split}&L^{1}_{V}\left({\bf f}\otimes x\right)=\left({\rm id}\otimes\chi\otimes{\rm id}\right)(12)\left({\rm id}\otimes\lambda^{1}_{P_{H}(V)}\right)\left(\nu\otimes{\rm id}\right)\left({\bf f}\otimes x\right)\\ &=\left({\rm id}\otimes\chi\otimes{\rm id}\right)(12)({\rm id}\otimes\lambda^{1}_{P_{H}(V)})\left({\bf f}\otimes 1\otimes x\right)\\ &=\left({\rm id}\otimes\chi\otimes{\rm id}\right)\left(\psi^{1}\otimes{\bf f}\otimes{\rm S}(\psi^{2})\otimes x\right)\\ &=\psi^{1}\otimes\chi({\bf f}\otimes{\rm S}(\psi^{2}))\otimes x=1\otimes{\bf f}\otimes x,\,\,\,{\rm and}\,\,\,\end{split}$
$\begin{split}&R^{1}_{V}\left({\bf f}\otimes x\right)=\left(\chi\otimes(12)\right)\left({\rm id}\otimes\rho^{1}_{P_{H}(V)}\right)\left(\nu\otimes{\rm id}\right)\left({\bf f}\otimes x\right)\\ &=(\chi\otimes(12))({\rm id}\otimes\rho^{1}_{P_{H}(V)})\left({\bf f}\otimes 1\otimes x\right)=(\chi\otimes(12))\left({\bf f}\otimes 1\otimes x^{0}\otimes x^{-1}\right)\\ &=\chi({\bf f}\otimes 1)\otimes x^{-1}\otimes x^{0}={\bf f}\otimes x^{-1}\otimes x^{0},\end{split}$
as claimed. ∎
Example 6.7.
Let $\mathfrak{g}$ be a finite dimensional restricted $p$-Lie algebra, $\mathfrak{h}\subset\mathfrak{g}$ a restricted $p$-Lie subalgebra, and $G,H$ the associated finite group schemes. Consider the finite tensor category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$. (See [G3] for examples of nontrivial cocycles.) By Theorem 6.6, we have ${\mathscr{C}}=\overline{{\rm Rep}(\mathfrak{h})_{k}}$.
For example, if $\mathfrak{h}$ is unipotent then $\mathscr{C}$} is a unipotent tensor category with unique simple object $\mathbf{1}:=\mathbf{F}(k)=\mathscr{O}(H)$, and
$P_{{\mathscr{C}}}(\mathbf{1})\cong\mathscr{O}(G)\otimes_{k}P_{H}(k)\cong\mathscr{O}(G)\otimes_{k}\mathscr{O}(H)$
is the free $\mathscr{O}(G)$-module of rank $|H|$ with trivial left $\mathscr{O}(H)$-cocation and right $\mathscr{O}(H)$-cocation ${\rm id}\otimes\Delta_{\psi}$. So, ${\mathscr{C}}\simeq{\rm Coh}(G,\omega)$ as abelian categories, but not necessarily as tensor categories (see [GS, Example 6.8]). ∎
6.3. The normal case
Assume $H$ is normal in $G$. Then $\GG/H=H\backslash G$ is a finite group scheme, and the quotient morphism $\pi:G\to G/H$ (2.17) is a group scheme morphism.
Recall (2.23) the maps $\rho^{\psi}_{U}$, $U\in{\rm Coh}(G/H)$.
Theorem 6.8.
The following hold:
  1. (1)
    The injective tensor functor $\pi^{*}:{\rm Coh}(G/H)\xrightarrow{1:1}{\rm Coh}(G)$ lifts to an injective tensor functor
    $\pi^{*}:{\rm Coh}(G/H)\xrightarrow{1:1}{\mathscr{C}},\,\,\,U\mapsto\left(\pi^{*}U,\left(\rho^{\psi}_{U}\right)_{21},\rho^{\psi}_{U}\right).$
  2. (2)
    Set $\mathbf{F}:=\mathbf{F}_{H}$. We have an equivalence of abelian categories
    ${\rm Coh}(G/H)\boxtimes{\rm Rep}(H)_{k}\xrightarrow{\simeq}{\mathscr{C}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V).$
Proof.
444 See [BG, Example 5.8] for a different proof.
(1) Follows from Theorem 2.16 (as $H$ is normal).
(2) Since for any $\bar{g}\in(G/H)(k)$, the object $\pi^{*}\delta_{\bar{g}}\in{\mathscr{C}}_{gH}$ is invertible, we have an equivalence of abelian categories
$\pi^{*}\delta_{\bar{g}}\otimes-:\overline{{\mathscr{C}}_{H}}\xrightarrow{\simeq}\overline{{\mathscr{C}}_{gH}},\,\,\,S\mapsto\pi^{*}\delta_{\bar{g}}\otimes S.$
Now since we have an equivalence
${\rm Coh}(G/H)_{\bar{1}}\boxtimes{\rm Rep}(H)_{k}\xrightarrow{\simeq}\overline{{\mathscr{C}}_{H}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V),$
it follows that for any $\bar{g}\in(G/H)(k)$, the functor
${\rm Coh}(G/H)_{\bar{g}}\boxtimes{\rm Rep}(H)_{k}\xrightarrow{\simeq}\overline{{\mathscr{C}}_{gH}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V),$
is an equivalence of abelian categories, which implies the statement. ∎
7. The center of ${\rm Coh}(G,\omega)$
Throughout this section, we fix $(G,\omega)$ 2.7. Let $\Delta\colon G\to G\times G$ be the diagonal map, $\mathbb{G}:=G\times G$, and $\mathbb{H}:=\Delta(G)$.
Let $\mathscr{Z}(G,\omega):=\mathscr{Z}({\rm Coh}(G,\omega))$ be the center of ${\rm Coh}(G,\omega)$. Recall that objects of $\mathscr{Z}(G,\omega)$ are pairs $(X,{\rm c})$, where $X\in{\rm Coh}(G,\omega)$ and
${\rm c}:(-\otimes X)\xrightarrow{\cong}(X\otimes-)$
is a natural isomorphism satisfying a certain property, usually known as a half-braiding (see, e.g., [DGNO, Section 7.13]). The center $\mathscr{Z}(G,\omega)$ is a finite nondegenerate braided tensor category (see, e.g., [DGNO, Section 8.6.3]).
Recall that there is a canonical equivalence of tensor categories
(7.1) $\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\left({\rm Coh}(G,\omega)\boxtimes{\rm Coh}\left(G,\omega\right)^{{}^{\rm rev}}\right)^{*}_{{\rm Coh}(G,\omega)}$
assigning to a pair $(X,{\rm c})$ the functor $X\otimes-\colon{\rm Coh}(G,\omega)\to{\rm Coh}(G,\omega)$, equipped with a module structure coming from $\rm c$} (see, e.g., [DGNO, Proposition 7.13.8]).
Let $p_{1},p_{2}:\mathbb{G}\twoheadrightarrow G$ be the obvious projection morphisms, and let
$\widetilde{\omega}:=p_{1}^{\sharp\otimes 3}(\omega)\cdot p_{2}^{\sharp\otimes 3}(\omega^{-1})\in Z^{3}(\mathbb{G},\mathbb{G}_{m}).$
Then there is a canonical equivalence of tensor categories
(7.2) ${\rm Coh}(\mathbb{G},\widetilde{\omega})^{*}_{{\mathscr{M}}(\mathbb{H},1)}\xrightarrow{\simeq}\left({\rm Coh}(G,\omega)\boxtimes{\rm Coh}\left(G,\omega\right)^{{}^{\rm rev}}\right)^{*}_{{\rm Coh}(G,\omega)}.$
Thus, (7.1)–(7.2) yield a canonical equivalence of tensor categories
(7.3) ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)\simeq\mathscr{Z}(G,\omega).$
Finally, recall that there is a canonical tensor equivalence
(7.4) ${\rm Coh}^{(G)}(G,\omega)\simeq\mathscr{Z}(G,\omega),$
where ${\rm Coh}^{(G)}(G,\omega)$ is the category of right $G$-equivariant sheaves on $(G,\omega)$ with respect to right conjugation (see Definition 2.12). Thus, (7.3)–(7.4) yield a canonical equivalence of tensor categories
(7.5) ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)\simeq{\rm Coh}^{(G)}(G,\omega).$
In this section, we study the tensor category $\mathscr{Z}(G,\omega)$ from two perspectives. First, we use Theorem 6.2, and the descriptions of $\mathscr{Z}(G,\omega)$ mentioned above, to study the abelian structure of $\mathscr{Z}(G,\omega)$. We then use (2.4) and [GNN] to describe $\mathscr{Z}(G,\omega)$ as a $G(k)$-equivariantization. Each of these descriptions provides a certain direct sum decomposition of $\mathscr{Z}(G,\omega)$, and we end the discussion by establishing a relation between the components coming from the two decompositions.
7.1. The structure of ${\mathscr{C}}\left(\mathbb{G},\tilde{\omega},\mathbb{H},1\right)$
Let ${\rm Y}:=\mathbb{G}/(\mathbb{H}\times\mathbb{H})$ with respect to the right action $\mu_{\mathbb{G}\times(\mathbb{H}\times\mathbb{H})}$ (4.1). Note that for any closed point ${\rm g}\in\mathbb{G}(k)$, we have $\xi_{{\rm g}}=\iota_{{\rm g}}^{\sharp\otimes 2}({\rm w}_{{\rm g}})$ (4.7).
Theorem 7.1.
Let ${\mathscr{C}}:={\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)$. The following hold:
  1. (1)
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, we have an equivalence of abelian categories
    $\mathbf{F}_{{\rm Z}}:{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}\xrightarrow{\simeq}\mathscr{C}_{{\rm Z}},\,\,\,V\mapsto\iota_{{\rm Z}*}\left(\mathscr{O}\left({\rm Z}\right)\otimes_{k}V,\lambda_{V}^{{\rm g}},\rho^{{\rm g}}_{V}\right).$
    In particular, we have a tensor equivalence
    $\mathbf{F}_{\mathbb{H}}:{\rm Rep}(\mathbb{H})_{k}\xrightarrow{\simeq}\mathscr{C}_{\mathbb{H}},\,\,\,V\mapsto\iota_{\mathbb{H}*}\left(\mathscr{O}\left(\mathbb{H}\right)\otimes_{k}V,\lambda_{V}^{1},\rho^{1}_{V}\right).$
  2. (2)
    There is a bijection between equivalence classes of pairs $({\rm Z},V)$, where ${\rm Z}\in{\rm Y}(k)$ is a closed point with representative ${\rm g}\in{\rm Z}(k)$, and $V\in{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$ is simple, and simple objects of $\mathscr{C}$}, assigning $({\rm Z},V)$ to $\mathbf{F}_{{\rm Z}}(V)$. Moreover, we have a direct sum decomposition of abelian categories
    ${\mathscr{C}}=\bigoplus_{{\rm Z}\in{\rm Y}(k)}\overline{{\mathscr{C}}_{{\rm Z}}},$
    and $\overline{{\mathscr{C}}_{\mathbb{H}}}\subset{\mathscr{C}}$ is a tensor subcategory.
  3. (3)
    For any $V\in{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$, we have $\mathbf{F}_{{\rm Z}}(V)^{*}\cong\mathbf{F}_{{\rm Z}^{-1}}(V^{*})$.
  4. (4)
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, and $V$ in ${\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$, we have
    ${\rm FPdim}\left(\mathbf{F}_{{\rm Z}}(V)\right)=\frac{|\mathbb{H}|}{|\mathbb{H}^{{\rm g}}|}{\rm dim}(V).$
  5. (5)
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, and simple $V\in{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$, we have
    $P_{{\mathscr{C}}}\left(\mathbf{F}_{{\rm Z}}(V)\right)\cong\left(\mathscr{O}(\mathbb{G}^{\circ})\otimes\mathscr{O}({\rm Z}(k))\otimes_{k}P_{(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})}(V),L^{{\rm g}}_{V},R^{{\rm g}}_{V}\right),\,\,\,{\rm and}\,\,\,$
    ${\rm FPdim}\left(P_{{\mathscr{C}}}\left(\mathbf{F}_{{\rm Z}}\left(V\right)\right)\right)=\frac{|\mathbb{G}^{\circ}||\mathbb{H}(k)|}{|\mathbb{H}^{\circ}||\mathbb{H}^{{\rm g}}(k)|}{\rm dim}\left(P_{(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})}\left(V\right)\right).$
  6. (6)
    For any ${\rm Z}\in{\rm Y}(k)$, we have ${\rm FPdim}\left(\overline{{\mathscr{C}}_{{\rm Z}}}\right)=|\mathbb{G}^{\circ}||{\rm Z}(k)|$.
Proof.
Follows immediately from Theorem 6.2. ∎
Corollary 7.2.
[…]
Proof.
Follows from Corollary 5.5. ∎
7.2. The structure of ${\rm Coh}^{(G)}(G,\omega)$
Consider the right conjugation action of $G$ on itself. Let $\rm C$} be the finite scheme of conjugacy orbits in $G$. Then for any closed point $C\in{\rm C}(k)$, $C\subset G$ is closed and $C(k)\subset G(k)$ is a conjugacy class. Fix a representative $g=g_{C}\in C(k)$, and let $G_{C}$ denote the centralizer of $g$ in $G$ (so $G_{C}(k)$ is the centralizer of $g$ in $G(k)$).
Note that the map
${\rm C}(k)\to{\rm Y}(k),\,\,\,C_{g}\mapsto{\rm Z}_{(g,1)},$
is bijective with inverse given by
${\rm Y}(k)\to{\rm C}(k),\,\,\,{\rm Z}_{(g_{1},g_{2})}\mapsto C_{g_{1}g_{2}^{-1}}.$
Also, for any $C\in{\rm C}(k)$ with representative $g\in C(k)$, we have
$\mathbb{H}^{(g,1)}=\mathbb{H}\cap(g,1)\mathbb{H}(g^{-1},1)=\Delta(G_{C}).$
For any $C\in{\rm C}(k)$ with representative $g\in C(k)$, let $\iota_{g}:G_{C}\hookrightarrow G$ be the inclusion morphism. Then we have the following lemma whose proof is similar to the proof of Lemma 4.1.
Lemma 7.3.
For any $C\in{\rm C}(k)$ with representative $g\in C(k)$, the element $\iota_{g}^{\sharp\otimes 2}(\omega_{g})$ (2.15) lies in $Z^{2}(G_{C},\mathbb{G}_{m})$. ∎
Now choose a cleaving map (2.29) $\mathfrak{c}_{g}:\mathscr{O}(G_{C})\xrightarrow{1:1}\mathscr{O}(G)$, and let
[…]
(see (2.32)). Consider also the split exact sequence of schemes
$1\to C^{\circ}\xrightarrow{i_{C^{\circ}}}C\mathrel{\mathop{\rightleftarrows}^{\pi_{C}}_{q_{C}}}C(k)\to 1$
induced from (2.3), and define the $\mathscr{O}(G)$-linear algebra maps
$\chi_{C}:={\rm id}\otimes q_{C}^{\sharp}:\mathscr{O}(G^{\circ})\otimes\mathscr{O}(C)\twoheadrightarrow\mathscr{O}(G^{\circ})\otimes\mathscr{O}(C(k)),\,\,\,{\rm and}$
$\nu_{C}:={\rm id}\otimes\pi_{C}^{\sharp}:\mathscr{O}(G^{\circ})\otimes\mathscr{O}(C(k))\xrightarrow{1:1}\mathscr{O}(G^{\circ})\otimes\mathscr{O}(C).$
Theorem 7.4.
Set $\mathscr{Z}:={\rm Coh}^{(G)}(G,\omega)$. The following hold:
  1. (1)
    For any $C\in{\rm C}(k)$ with representative $g\in C(k)$, we have an equivalence of abelian categories
    $\mathbf{F}_{C}:{\rm Rep}_{k}(G_{C},\omega_{g})\xrightarrow{\simeq}\mathscr{Z}_{C},\,\,\,V\mapsto\iota_{C*}\left(\mathscr{O}(C)\otimes_{k}V,\rho_{V}^{g}\right),$
    where $\rho_{V}^{g}:\mathscr{O}(C)\otimes_{k}V\to\mathscr{O}(C)\otimes_{k}V\otimes\mathscr{O}(G)$ is given by
    $\rho_{V}^{g}(f\otimes v)=\left(\mathfrak{j}_{g}^{-1}\right)^{\sharp}\alpha_{g}\left(\mathfrak{j}_{g}^{\sharp}(f)_{1}\mathfrak{c}_{g}\left(v^{1}\right)_{1}\right)\otimes v^{0}\otimes\mathfrak{j}_{g}^{\sharp}(f)_{2}\mathfrak{c}_{g}\left(v^{1}\right)_{2}.$
[…]
    In particular, $\mathbf{F}_{1}:{\rm Rep}_{k}(G)\xrightarrow{\simeq}\mathscr{Z}_{1}\hookrightarrow\mathscr{Z}$ coincides with the canonical embedding of braided tensor categories.
  2. (2)
    There is a bijection between equivalence classes of pairs $(C,V)$, where $C\in{\rm C}(k)$ is a closed point with representative $g\in C(k)$, and $V$ in ${\rm Rep}_{k}(G_{C},\omega_{g})$ is simple, and simple objects of $\mathscr{Z}$, assigning $(C,V)$ to $\mathbf{F}_{C}(V)$. Moreover, we have a direct sum decomposition of abelian categories
    $\mathscr{Z}=\bigoplus_{C\in{\rm C}(k)}\overline{\mathscr{Z}_{C}},$
    and $\overline{\mathscr{Z}_{1}}\subset\mathscr{Z}$ is a tensor subcategory.
  3. (3)
    For any $V\in{\rm Rep}_{k}(G_{C},\omega_{g})$, we have $\mathbf{F}_{C}(V)^{*}\cong\mathbf{F}_{C^{-1}}(V^{*})$.
  4. (4)
    For any $V\in{\rm Rep}_{k}(G_{C},\omega_{g})$, we have
    ${\rm FPdim}\left(\mathbf{F}_{C}(V)\right)=\frac{|G|}{|G_{C}|}{\rm dim}(V)=|C|{\rm dim}(V).$
  5. (5)
    For any simple $V\in{\rm Rep}_{k}(G_{C},\omega_{g})$, we have
    $P_{\mathscr{Z}}\left(\mathbf{F}_{C}(V)\right)\cong\left(\mathscr{O}(G^{\circ})\otimes\mathscr{O}(C(k))\otimes_{k}P_{(G_{C},\omega_{g})}(V),R_{V}^{g}\right),$
    where $\mathscr{O}(G)$ acts diagonally,
    $R_{V}^{g}:=r^{g}\cdot\left(\chi_{C}\otimes{\rm id}^{\otimes 2}\right)\left({\rm id}_{\mathscr{O}(G^{\circ})}\otimes\rho_{P_{(G_{C},\omega_{g})}(V)}^{g}\right)\left(\nu_{C}\otimes{\rm id}\right),$
    and $r^{g}:=(\iota_{G^{\circ}}^{\sharp}\otimes\iota_{C}^{\sharp}\otimes{\rm id})(\omega)$. In particular,
    ${\rm FPdim}\left(P_{\mathscr{Z}}\left(\mathbf{F}_{C}(V)\right)\right)=\frac{|G|}{|G_{C}(k)|}{\rm dim}\left(P_{(G_{C},\omega_{g})}\left(V\right)\right).$
  6. (6)
    For any $C\in{\rm C}(k)$, we have ${\rm FPdim}\left(\overline{\mathscr{Z}_{C}}\right)=\frac{|G|^{2}}{|G_{C}(k)|}$. ∎
Proof.
Using the preceding remarks, it is straightforward to verify that Theorem 7.1 translates to the theorem via the equivalence (7.5). ∎
Example 7.5.
Assume $G$ is connected (e.g., $G$ is the finite group scheme associated to a finite dimensional restricted $p$-Lie algebra $\mathfrak{g}$). Assume that $\omega\in Z^{3}(G,\mathbb{G}_{m})$, and consider the finite braided tensor category $\mathscr{Z}:=\mathscr{Z}\left(G,\omega\right)$. By Theorem 7.4, we have $\mathscr{Z}\left(G,\omega\right)=\overline{{\rm Rep}_{k}(G)}$, and $\rho^{1}_{V}=\rho_{V}:V\to V\otimes\mathscr{O}(G)$ for any $(V,\rho_{V})$ in ${\rm Rep}_{k}(G)$. Moreover, for any simple $(V,\rho_{V})\in{\rm Rep}_{k}(G)$, we have
$P_{\mathscr{Z}}\left(\mathbf{F}(V)\right)\cong\left(\mathscr{O}(G)\otimes_{k}P_{G}(V),R_{V}^{1}\right),$
where $\mathscr{O}(G)$ acts on the first factor, and
$R_{V}^{1}:=r^{1}\cdot\left(\chi_{1}\otimes{\rm id}^{\otimes 2}\right)\left({\rm id}_{\mathscr{O}(G)}\otimes\rho^{1}_{P_{G}(V)}\right)\left(\nu_{1}\otimes{\rm id}\right).$
Now since
$\nu_{1}:\mathscr{O}(G)\xrightarrow{1:1}\mathscr{O}(G)\otimes\mathscr{O}(G),\,\,\,f\mapsto f\otimes 1,$
$\chi_{1}:\mathscr{O}(G)\otimes\mathscr{O}(G)\twoheadrightarrow\mathscr{O}(G),\,\,\,f\otimes f^{\prime}\mapsto f\varepsilon(f^{\prime}),$
$\mathfrak{c}_{1}={\rm id},\,\,\,{\rm and}\,\,\,\alpha_{1}=\varepsilon,$
it follows that $R_{V}^{1}={\rm id}_{\mathscr{O}(G)}\otimes\rho_{P_{G}(V)}$. Thus,
$P_{\mathscr{Z}}\left(\mathbf{F}(V)\right)\cong\left(\mathscr{O}(G)\otimes_{k}P_{G}(V),{\rm id}\otimes\rho_{P_{G}(V)}\right),$
and $r^{1}=1$. ∎
7.3. Short exact sequence of centers
Recall the split exact sequence of group schemes
$1\to G^{\circ}\xrightarrow{i}G\mathrel{\mathop{\rightleftarrows}^{\pi}_{q}}G(k)\to 1$
(see (2.3)), and set
(7.7) $\omega^{\circ}:=\iota^{\sharp\otimes 3}(\omega)\in Z^{3}(G^{\circ},\mathbb{G}_{m}),\,\,\,\,\omega(k):=q^{\sharp\otimes 3}(\omega)\in Z^{3}(G(k),\mathbb{G}_{m}).$
Note that $\pi^{\sharp\otimes 3}(\omega(k))=\omega$, and (2.3) induces the tensor functors
$i_{*}:{\rm Coh}(G^{\circ},\omega^{\circ})\xrightarrow{1:1}{\rm Coh}(G,\omega),\,\,\,i^{*}:{\rm Coh}(G,\omega)\twoheadrightarrow{\rm Coh}(G^{\circ},\omega^{\circ}),$
$\pi_{*}:{\rm Coh}(G,\omega)\twoheadrightarrow{\rm Coh}(G(k),\omega(k)),\,\,\,q_{*}:{\rm Coh}(G(k),\omega(k))\xrightarrow{1:1}{\rm Coh}(G,\omega).$
(See Proposition 2.8.)
Theorem 7.6.
The following hold:
  1. (1)
    The functor $\pi_{*}$ lifts (using $q$) to a surjective quasi-tensor functor
    $\pi_{*}:{\rm Coh}^{(G)}(G,\omega)\twoheadrightarrow{\rm Coh}^{(G(k))}(G(k),\omega(k)),$
    and if $\omega=1$, then it is a tensor functor.
  2. (2)
    The functor $q_{*}$ lifts (using $\pi$) to an injective tensor functor
    $q_{*}:{\rm Coh}^{(G(k))}(G(k),\omega(k))\xrightarrow{1:1}{\rm Coh}^{(G)}(G,\omega).$
  3. (3)
    $\pi_{*}q_{*}={\rm id}$ (as abelian functors).
  4. (4)
    The functor $i^{*}$ lifts to a surjective quasi-tensor functor
    $i^{*}:{\rm Coh}^{(G)}(G,\omega)\twoheadrightarrow{\rm Coh}^{(G^{\circ})}(G^{\circ},\omega^{\circ}),$
    and if $\omega=1$, then it is a tensor functor.
  5. (5)
    $i^{*}q_{*}:{\rm Coh}^{(G(k))}(G(k),\omega(k))\to{\rm Vec}=\langle\mathbf{1}\rangle\subset{\rm Coh}^{(G^{\circ})}(G^{\circ},\omega^{\circ})$ is the forgetful functor (as abelian functors).
  6. (6)
    The identity functor ${\rm Coh}(G,\omega)\to{\rm Coh}(G,\omega)$ lifts (using $i$) to a surjective tensor functor
    ${\rm Coh}^{(G)}(G,\omega)\twoheadrightarrow{\rm Coh}^{(G^{\circ})}(G,\omega).$
Proof.
(1) Take $(V,\rho)$ in ${\rm Coh}^{(G)}(G,\omega)$. Namely, $V$ is an $\mathscr{O}(G)$-module and $\rho:V\to V\otimes\mathscr{O}(G)$ is an $\mathscr{O}(G,\omega)$-coaction with respect to right conjugation (see Definition 2.12). By definition, $\pi_{*}V=V$ is an $\mathscr{O}(G(k))$-module via $\pi^{\sharp}:\mathscr{O}(G(k))\xrightarrow{1:1}\mathscr{O}(G)$. Thus, the $(\mathscr{O}(G(k)),\omega(k))$-coaction
$({\rm id}\otimes q^{\sharp})\rho:\pi_{*}V\to\pi_{*}V\otimes\mathscr{O}(G(k))$
endows $\pi_{*}V$ with a structure of an object in ${\rm Coh}^{(G(k))}(G(k),\omega(k))$.
(2) Take $(V,\rho)$ in ${\rm Coh}^{(G(k))}(G(k),\omega(k))$. Namely, $V$ is an $\mathscr{O}(G(k))$-module and $\rho:V\to V\otimes\mathscr{O}(G(k))$ is an $(\mathscr{O}(G(k)),\omega(k))$-coaction with respect to right conjugation. Now by definition, $q_{*}V=V$ is an $\mathscr{O}(G)$-module via $q^{\sharp}:\mathscr{O}(G)\twoheadrightarrow\mathscr{O}(G(k))$. Thus, the $(\mathscr{O}(G),\omega)$-coaction
$({\rm id}\otimes\pi^{\sharp})\rho:q_{*}V\to q_{*}V\otimes\mathscr{O}(G)$
endows $q_{*}V$ with a structure of an object in ${\rm Coh}^{(G)}(G,\omega)$.
(3) Follows from $\pi_{*}q_{*}=(\pi q)_{*}={\rm id}$.
(4) Take $(V,\rho)$ in ${\rm Coh}^{(G)}(G,\omega)$. By Proposition 2.8, the $\mathscr{O}(G^{\circ})$-module $i^{*}V=\mathscr{O}\left(G^{\circ}\right)\otimes_{\mathscr{O}\left(G\right)}V$ is equipped with the $(\mathscr{O}(G),\omega)$-coaction given by $\mu_{G^{\circ}\times G}^{\sharp}\bar{\otimes}\rho$ (where $\mu_{G^{\circ}\times G}:G^{\circ}\times G\to G^{\circ}$ is the right conjugation action of $G$ on $G^{\circ}$). Thus, the map
$\left({\rm id}\otimes{\rm id}\otimes i^{\sharp}\right)\left(\mu_{G^{\circ}\times G}^{\sharp}\bar{\otimes}\rho\right),$
endows $i^{*}V$ with a structure of an object in ${\rm Coh}^{(G^{\circ})}(G^{\circ},\omega^{\circ})$.
(5) Take $(V,\rho)$ in ${\rm Coh}^{(G(k))}(G(k),\omega(k))$. By (2),
$q_{*}(V,\rho)=(q_{*}V,({\rm id}\otimes\pi^{\sharp})\rho)\in{\rm Coh}^{(G)}(G,\omega),$
where $q_{*}V=V$ is an $\mathscr{O}(G)$-module via $q^{\sharp}$. Thus by (4),
$i^{*}q_{*}(V,\rho)=i^{*}\left(q_{*}V,\left({\rm id}\otimes\pi^{\sharp}\right)\rho\right)$
$=$ $\left(\mathscr{O}\left(G^{\circ}\right)\otimes_{\mathscr{O}\left(G\right)}q_{*}V,\left({\rm id}\otimes{\rm id}\otimes i^{\sharp}\right)\left(\mu_{G^{\circ}\times G}^{\sharp}\bar{\otimes}\left({\rm id}\otimes\pi^{\sharp}\right)\rho\right)\right)$
$=$ $\left(\mathscr{O}\left(G^{\circ}\right)\otimes_{\mathscr{O}\left(G\right)}q_{*}V,\mu_{G^{\circ}\times G^{\circ}}^{\sharp}\bar{\otimes}V_{{\rm tr}}\right)$
(where $\mu_{G^{\circ}\times G^{\circ}}:G^{\circ}\times G^{\circ}\to G^{\circ}$ is the right conjugation action of $G^{\circ}$ on itself). In other words, $i^{*}q_{*}$ sends $(V,\rho)$ to the direct sum of $\dim(V)$ copies of the identity object $\left(\mathscr{O}(G^{\circ}),\mu_{G^{\circ}\times G^{\circ}}^{\sharp}\right)\in{\rm Coh}^{(G^{\circ})}(G^{\circ},\omega^{\circ})$.
(6) Similar. ∎
7.4. $\mathscr{Z}(G,\omega)$ as $G(k)$-equivariantization
Set $\mathscr{D}:={\rm Coh}(G,\omega)$, and $\mathscr{D}^{\circ}:={\rm Coh}(G^{\circ},\omega^{\circ})$. By (2.4), we have equivalences
$\mathscr{D}\simeq\mathscr{D}^{\circ}\rtimes G(k)\simeq\bigoplus_{a\in G(k)}\mathscr{D}^{\circ}\boxtimes a$
of tensor categories, where the associativity constraint on the right hand side category is given by $\omega$ in the obvious way.
Let $\mathscr{M}$} be any $\mathscr{D}^{\circ}$-bimodule category. Recall [GNN] that the relative center $\mathscr{Z}_{\mathscr{D}^{\circ}}\left({\mathscr{M}}\right)$ is the abelian category whose objects are pairs $(M,{\rm c})$, where $M\in{\mathscr{M}}$ and
(7.8) ${\rm c}=\{{\rm c}_{X}:X\otimes M\xrightarrow{\cong}M\otimes X\mid X\in\mathscr{D}^{\circ}\}$
is a natural family of isomorphisms satisfying some compatibility conditions. In particular, the relative center $\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)$ is a finite $G(k)$-crossed braided tensor category [GNN, Theorem 3.3]. The $G(k)$-grading on $\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)$ is given by
$\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)=\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right),$
and the action of $G(k)$ on $\mathscr{Z}_{\mathscr{D}^{\circ}}(\mathscr{D})$, $h\mapsto\widetilde{T}_{h}$, is induced from the action of $G(k)$ on $\mathscr{D}^{\circ}$, $h\mapsto T_{h}$, in the following way. For any $h\in G(k)$, $X\in\mathscr{D}^{\circ}$, and $(Y\boxtimes a,{\rm c})\in\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)$, we have an isomorphism
(7.9) $\widetilde{{\rm c}}_{X}:=(T_{h}\otimes T_{h}){\rm c}_{T_{h^{-1}}(X)}:X\otimes T_{h}(Y)\xrightarrow{\cong}T_{h}(Y)\otimes T_{hah^{-1}}(X).$
Set
[…]
Then $\widetilde{T}_{h}$ maps $\mathscr{Z}_{\mathscr{D}^{\circ}}(\mathscr{D}^{\circ}\boxtimes a)$ to $\mathscr{Z}_{\mathscr{D}^{\circ}}(\mathscr{D}^{\circ}\boxtimes{hah^{-1}})$.
Note that $\mathscr{Z}_{\mathscr{D}^{\circ}}(\mathscr{D})\simeq{\rm Coh}^{(G^{\circ})}(G,\omega)$ as tensor categories, and the obvious forgetful tensor functor
(7.10) $\mathscr{Z}(G,\omega)=\mathscr{Z}_{\mathscr{D}}\left(\mathscr{D}\right)\to\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right),\,\,\,(X,{\rm c})\mapsto(X,{\rm c}_{\mid\mathscr{D}^{\circ}})$
coincides with the surjective tensor functor given in Theorem 7.6(6).
By [GNN, Theorem 3.5], there is an equivalence of tensor categories
(7.11) $F:\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}=\left(\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)\right)^{G(k)}.$
For any $C\in{\rm C}(k)$, set
$\mathscr{E}_{C}:=\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}\subset\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}.$
By (7.11), $\mathscr{E}_{C}$ is a Serre subcategory of $\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}$, and we have a tensor equivalence
(7.12) $F:\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\bigoplus_{C\in{\rm C}(k)}\mathscr{E}_{C}.$
Theorem 7.7.
For any $C\in{\rm C}(k)$, the functor $F$ (7.12) restricts to an equivalence of abelian categories
$F_{C}:\overline{\mathscr{Z}(G,\omega)_{C}}\xrightarrow{\simeq}\mathscr{E}_{C}=\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}.$
In particular, $F$ restricts to an equivalence of tensor categories
$F_{1}:\overline{{\rm Rep}_{k}(G)}\xrightarrow{\simeq}\mathscr{Z}\left(G^{\circ},\omega^{\circ}\right)^{G(k)}.$
Proof.
Fix $C\in{\rm C}(k)$ with representative $g\in C(k)$. To see that $F\left(\overline{\mathscr{Z}(G,\omega)_{C}}\right)\subset\mathscr{E}_{C}$ it is enough to show that $F\left(\mathscr{Z}(G,\omega)_{C}\right)\subset\mathscr{E}_{C}$ (by the discussion above). To this end, it is enough to show that for any simple ${\rm V}\in\mathscr{Z}(G,\omega)_{C}$, the simple $G(k)$-equivariant object $F({\rm V})\in\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)$ is supported on $C$.
So, let ${\rm V}\in\mathscr{Z}(G,\omega)_{C}$ be simple. By Theorem 7.4, there exists a unique simple $V$ in ${\rm Rep}_{k}(G_{C},\omega_{g})$ such that ${\rm V}={\bf F}_{C}(V)=\mathscr{O}(C)\otimes_{k}V$. It follows that the forgetful image of $F({\rm V})$ in $\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)$ (7.10) lies in $\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)$, so by the proof of [GNN, Theorem 3.5], $F({\rm V})$ is supported on $C$, as claimed. ∎